### v1

# “Sharing, Not Showing Off”: How BeReal Approaches Authentic Self-Presentation on Social Media Through Its Design

Abstract.
Adolescents are particularly vulnerable to the pressures created by social media, such as heightened self-consciousness and the need for extensive self-presentation. In this study, we investigate how BeReal, a social media platform designed to counter some of these pressures, influences adolescents’ self-presentation behaviors. We interviewed 29 users aged 13-18 to understand their experiences with BeReal. We found that BeReal’s design focuses on spontaneous sharing, including randomly timed daily notifications and reciprocal posting, discourages staged posts, encourages careful curation of the audience, and reduces pressure on self-presentation. The space created by BeReal offers benefits such as validating an unfiltered life and reframing social comparison, but its approach to self-presentation is sometimes perceived as limited or unappealing and, at times, even toxic. Drawing on this empirical data, we distill a set of design guidelines for creating platforms that support authentic self-presentation online, such as scaffolding reciprocity and expanding beyond spontaneous photo-sharing to allow users to more accurately and comfortably portray themselves.
1. Introduction
The number of people who report having no one to discuss difficult subjects with has tripled in recent decades, and the average person’s social network has shrunk by one-third (McPherson et al., 2006). A significant increase in social isolation has occurred, with nearly 25% of Americans in 2004 reporting they had no confidant, up from 10% in 1985 (McPherson et al., 2006). This is considering the profound importance of social relationships on our physical and mental health, longevity, and happiness (Cacioppo and Cacioppo, 2014; House et al., 1988; Cohen, 2004). The Study of Adult Development examined participants over 75 years and indicates that good relationships keep us happier and healthier (Waldinger and Schulz, 2016). A 2010 review of over 300,000 participants found that weak social ties are as detrimental to health as alcoholism (Holt-Lunstad et al., 2010). For adolescents, this intrinsic need for trusted interpersonal connection, particularly with peers, is even more pronounced than with adults (Erikson, 1994a; Davis, 2012). And with the widespread adoption of social media, the nature of these social interactions has transformed: 95% of adolescents ages 13-17 engage actively with friends through social media platforms (noa, 2023). As these peer relationships increasingly have an online dimension, it is crucial to assess how technology supports teens’ relationship-building.
Social Penetration Theory (Carpenter and Greene, 2015a) explains that relationships develop through reciprocal and authentic self-disclosure, which progresses in depth over time. However, social media often offers “empty calories of social interaction” (Prinstein, 2023), which might seem to fulfill adolescents’ developmental needs, but actually lack the substantive qualities necessary for true social benefits. The use of these platforms can amplify inherent developmental vulnerabilities of adolescents, such as increased sensitivity to social feedback and emotional reactivity due to ongoing brain maturation (Somerville, 2013a; Steinberg, 2005; Falk et al., 2014; Peake et al., 2013; Somerville, 2013b), which can further detract them from forming meaningful and supportive relationships. Nonetheless, a majority of teens use social media and regularly engage in extensive self-presentation through manicured content and the anxiety and distorted perception of reality that it can provoke (Aacap, [n. d.]).
BeReal is a social media platform created to combat the difficulty of cultivating relationships online due to the sparsity of authentic interactions. BeReal has a specific focus: “everyday get a notification to share a genuine glimpse into your real life with the people you care most about” which they believe enables users to “show your friends who you really are” (noa, [n. d.]). Though it was launched in 2019, long before the surgeon general’s advisory, its design and mission align with the surgeon general’s recommendation that tech companies “seek to maximize the potential benefits and avoid design features that attempt to maximize time, attention, and engagement” (noa, 2023).
On BeReal, users receive a notification informing them that they have a brief window of time to capture and share photos of their current activity and appearance, making strategic self-presentation nearly impossible. Consequently, the content shared on BeReal is more likely to be mundane or monotonous than content shared on other platforms. Yet, the app has become increasingly popular since early 2022, especially with youth (cycles and Text, [n. d.]). According to a 2023 Pew Research Report, BeReal usage among teens is comparable to that of Reddit or Twitch (Anderson et al., 2023). But regardless of the fate of the BeReal platform, it serves as a useful case study in designing to encourage authentic self-presentation. Self-presentation can build close relationships (when it is genuine) or provoke anxiety (when it is artificial). Thus, there is great value in creating spaces that encourage the former.
To the best of our knowledge, BeReal is the first major social media platform that explicitly purports to encourage authentic self-presentation. We sought to understand:
RQ1: How, if at all, does the design of BeReal influence users’ self-presentation on the platform?
RQ2: How do teens feel about self-presentation on BeReal?
To answer these questions, we conducted semi-structured interviews with 29 BeReal users aged 13 to 18, asking about their general experiences with BeReal. We found that BeReal’s features and affordances are effective in encouraging users to share casual, authentic snapshots of their daily lives. Features such as randomly timed daily notifications encourage authentic self-presentation via unstaged, unfiltered photo-sharing, and design elements like reciprocal posting encourage users to curate their audience, thereby making their self-presentation process less intentional or controlled. The participants saw many benefits of such a design. They especially appreciated that BeReal nudged users to collectively share more casual and less curated content, a change they said they sought on social media. However, some participants also felt that their ability to present themselves accurately and intentionally was too limited, and they explained that unfiltered photos taken at unplanned times are just one way of giving others insight into their lives. Others said that BeReal’s explicit prioritization of “authenticity” in self-presentation sometimes led to toxic competition and judgment about which users are truly being authentic.
In this paper, we provide empirical evidence of BeReal’s success in encouraging adolescent users to present themselves authentically. This serves as a proof of concept that design can significantly influence self-presentation. We also document the limitations of BeReal’s approach and outline design recommendations for creating a platform that encourages developmentally supportive online self-presentation for adolescents, such as scaffolding reciprocal sharing and supporting self-presentations that actors find to be accurate portrayals of themselves. Based on our findings, we introduce the Social Media Self-Presentation Matrix (SMSM). Grounded in theories on self-presentation goals and processes (Schlenker, 2003), this matrix presents ways that social media design can support different self-presentation strategies. It does so through four facets of self-presentation, spanning axes from automatic to controlled and from self-enhancing to self-expressive. Using the matrix, we highlight an untapped design opportunity to support controlled and self-expressive self-presentation, which could offer an alternative way for social media platforms to achieve the goals initially intended by BeReal. Further, we propose a design agenda to support better relationship-building online based on authenticity for teens on social media. We hope this work will encourage designers and researchers to more comprehensively support authentic exchanges in teen peer-relationship-building on social media, addressing a developmental need that is currently underrepresented.
2. Related Work
2.1. Adolescents and the Developmental Significance of Authentic Self-Disclosure
Forming close, trusting relationships is fundamental to people’s sense of meaning, happiness, and long-term well-being (Argyle, 2013). For adolescents, this intrinsic need for trusted interpersonal connections, particularly with peers, is even more pronounced than it is for adults (Erikson, 1994b; Davis, 2012). As described by the Social Penetration Theory (SPT) (Carpenter and Greene, 2015b), interpersonal relationships develop from initial, surface-level interactions to deeper, more intimate connections through a gradual process of self-disclosure. The breadth and depth of these disclosures increase as the relationship strengthens and matures. Also central to this progression is the norm of reciprocity, which posits that mutual exchange is fundamental to these evolving relationships.
For self-disclosure to effectively enhance intimacy and trust in relationships, it must be authentic. Authenticity—the characteristic of being in accordance with one’s internal values than with external contingencies (Peets and Hodges, 2018)—ensures that the information shared truly represents the individual’s self, fostering genuine connections and understanding between people (Laurenceau et al., 1998; Peets and Hodges, 2018). Moreover, expressing their authentic selves and receiving acknowledgment and affirmation of these selves from peers is crucial for individuals, particularly adolescents, to achieve a clear and coherent self-concept (Gardner, 2013). The formation of a coherent identity, rooted in an authentic self that is recognized and validated by one’s social circle, is fundamental to adolescent development (Erikson, 1994a). Research has also shown that authenticity in self-disclosure is associated with life satisfaction (Bailey et al., 2020), particularly among girls, linking authenticity to self-esteem and overall well-being (Impett et al., 2008). Moreover, a higher degree of authenticity in self-disclosing communication is connected to enhanced self-perceptions, reduced feelings of isolation in school, and greater satisfaction in relationships among adolescents (Peets and Hodges, 2018).
In recent years, particularly exacerbated by the COVID-19 pandemic and subsequent physical distancing measures, adolescents have increasingly turned to social media for social connection (Hamilton et al., 2022; Pitt et al., 2021). As of 2022, nearly all teenagers aged 13-17 (95%) own a smartphone or access the internet (97%), with nearly half (46%) using the internet ‘almost constantly’ (Vogels et al., 2022). Previous studies have identified that the primary motivation behind teens’ use of digital media is to maintain connections with existing friends (Ito, 2013; Boyd, 2008a; Davis, 2012). Given the critical role of online platforms in adolescent lives and the value placed on authenticity in their relationships, it is understandable that modern teenagers actively seek exchanges of authentic self-disclosure on social media and broader media contexts (Darr and Doss, 2022; Stephanie Rivas-Lara Becky Pham Jasmine Baten Annie Meyers Yalda T. Uhls, 2022). Further, the significantly reduced social cues on online platforms heighten uncertainty, making self-disclosure especially crucial for (adolescent) users to understand each other and strengthen their connections (Schlosser, 2020; Tidwell and Walther, 2002). This underscores the importance of examining how technology can support or undermine the development and maintenance of these relationships, with a particular focus on the reciprocal exchange of authentic self-disclosure throughout the stages of social penetration.
2.2. Authentic Self-Disclosure and Self-Presentation
Given that self-disclosure, especially in online contexts, involves extensive interpersonal contexts is social behavior that is likened a performance (Goffman, 2016) that symbolically communicates information about self to others (Blumer, 1986), self-presentation readily happens. Self-disclosure is defined as “verbally communicating personal information about the self to another person” (Forgas, 2011) (p. 449), and reflects communicating a factual representation of oneself, regardless of its effect on one’s public self-image. In contrast, self-presentation is defined as ‘‘the goal-directed activity of controlling information to influence the impressions formed by an audience about the self” (Schlenker and Wowra, 2003) (p. 871). Thus, self-disclosure typically aligns with authenticity, whereas self-presentation may suggest a lack of authenticity. However, these two are not mutually exclusive (Johnson, 1981); revealing true aspects of oneself does not necessarily equate to self-disclosure, as such revelations can also support a favorable public persona (Schlosser, 2020). Conversely, self-presentation should not be dismissed as “just superficial, deceptive, or manipulative” (Schlenker, 2003). In summary, we adopt the view in this paper that self-presentation is a strategic form of self-disclosure that predominates on social media, where users, acting as performers, anticipate positive outcomes from their disclosures. However, this form of self-presentation often conflicts with authenticity unless the self-presentation itself is authentic.
Although self-presentation is often perceived as strategic gamesmanship, this view does not capture its full complexity (Schlenker, 2003). In reality, self-presentation can be deeply “authentic,” as in, it typically involves a portrayal that is somewhat enhanced and idealized, yet genuinely believed by the person presenting it (Brown, 2014; Greenwald and Breckler, 1985; Williams and Gilovich, 2008). Sometimes the goal of self-presentation is accurate self-portrayal (Baumeister, 1982; Cheek and Hogan, 1983; Leary, 2019; Schlenker, 1980; Schlenker and Pontari, 2000). The aim could be to ensure that others perceive one in a manner that secures the respect and treatment aligned with one’s identity, to receive affirming feedback that addresses personal uncertainties, or to embody the principle that “honesty is the best policy,” thereby feeling authentic while reducing the risks associated with deception. Furthermore, intentional self-packaging does not necessarily indicate inauthenticity; akin to how a skilled lecturer carefully crafts their presentation to effectively convey genuine and significant information to their audience—contrasting starkly with a disorganized lecturer whose lack of preparation may seem insincere and self-absorbed—the careful arrangement of personal information can reflect an authentic depiction of oneself (as seen by the individual), not merely a false facade (Schlenker and Pontari, 2000). In other words, presenting accurate information requires as much skill in self-presentation as presenting inaccurate information if it is to achieve the intended effect on the audience.
Authentic self-presentation can manifest through either automatic or controlled cognitive processes. Automatic processes are characterized by 1) occurring without conscious awareness, 2) requiring minimal cognitive effort, 3) being autonomous and not requiring conscious monitoring, and 4) being initiated involuntarily without outside triggers (Bargh, 1996). In such instances, self-presentation scripts may be triggered automatically by specific aspects of the audience or the context, with the individuals often unaware of how deeply they are influenced by the social environment and their interpersonal objectives (Jones, 1990; Schlenker, 1980, 1985; Schlenker and Wowra, 2003; Tetlock and Manstead, 1985). Automatic self-presentation is most common in familiar environments among close acquaintances or friends, where individuals feel a sense of security and positive regard (e.g. while relaxing at home with closer friends) (Tyler, 2012). Self-presentation is also more likely to be automatic when it aligns with one’s established self-image and personality traits. Conversely, self-presentation often involves controlled processes in situations where the stakes or the audience are significant or where the individual is unsure about the impression they wish to project (Schlenker et al., 1996; Schlenker and Pontari, 2000). Under such conditions, individuals tend to deliberately focus on the impression they are making, meticulously planning and rehearsing their behavior (Tice et al., 1995; Schlenker et al., 1996). When required to deviate from their automatic mode of self-presentation, individuals may need to exert more cognitive effort, which can detract from their performance or quality of life (Vohs et al., 2005; Schlenker, 2003).
2.3. Challenges with Authentic Self-Presentation Online
While authentic self-presentation is highly valued, particularly among teenagers, it can be particularly challenging to achieve on social media. One characteristic of computer-mediated communication (CMC) that makes it more conducive to self-presentation than face-to-face (F2F) interactions is its asynchronicity (Schlosser, 2020). In contrast to the often messy, unscripted, and spontaneous nature of F2F communication, social media allows for more controlled exchanges due to the ability to reflect, edit, and revise messages (turkle2016reclaiming). Though self-presentation can also be spontaneous (Schlenker and Wowra, 2003), self-presentation is a complex process, reflecting a dynamic interplay between the self and the audience within a specific social context (Schlenker, 2003). It involves not only expressing oneself but also responding to situational pressures and conforming to the identity expectations of significant others (Schlenker et al., 1996). It embodies aspects of an individual’s self-concept, personality, social roles, and perceptions of audience preferences. Consequently, people often leverage the asynchronicity of CMC to craft their self-presentation online (vazire2004perceptions) carefully.
Online communication also differs from F2F interactions because it typically involves broadcasting messages to many people simultaneously rather than engaging in one-on-one exchanges (Schlosser, 2020). This broadcasting nature of social media interacts with adolescent characteristics in ways that significantly complicate self-presentation. Adolescents typically experience a heightened sense of self-consciousness relative to adults or younger children, leading to feelings of being under constant scrutiny, a phenomenon referred to as the imaginary audience (Elkind, 1967). This feeling is amplified in online environments where audiences are typically larger and less transparent than in offline scenarios (Zheng et al., 2019). Moreover, adolescents’ imagined audience on social media, defined as “the mental conceptualization of the people with whom [they] are communicating“ (Litt, 2012), encompasses individuals from diverse contexts such as family and school, mirroring the complex layers of audiences they engage with on these platforms. This sense of context collapse (Boyd, 2008b) prompts adolescents to employ self-presentation strategies in online environments to manage the tensions that arise (Marwick and Boyd, 2011; Boyd, 2008a; Tao and Ellison, 2023; Vitak, 2012). These strategies often involve disengaging with authentic self-presentation and instead sharing the “lowest common denominator” (Hogan, 2010): content that is universally deemed appropriate by all groups within the imagined audience (Hogan, 2010; Ullman, 1987). Consequently, relationship-building processes are also affected, as the breadth, depth, and reciprocity of authentic self-disclosure become restricted, and users disclose only superficial aspects of themselves (Hogan, 2010) or disengage entirely from sharing (Hargittai and Walejko, 2008).
The third distinguishing factor of many social media sites from F2F communication is the publicly visible and quantifiable nature of audience feedback, such as ‘likes’ and comments. Adolescents’ heightened self-consciousness often drives them to present themselves favorably online, striving to appear interesting, well-liked, and attractive to maintain their social standing (Yau and Reich, 2019; Haimson et al., 2021; Hamilton et al., 2022). They strive to expand their friend networks (since network size can suggest social embeddedness (Yau and Reich, 2019)), understand the social norms of the platforms (Naegele and Goffman, 1956), and create favorable profiles (often with assistance from close friends) to achieve peer acceptance (Yau and Reich, 2019). This leads to a tension between authenticity and self-presentation online known as the online authenticity paradox (Haimson et al., 2021): people aim for authenticity in their personal life and online presentation, but doing so necessitates sharing negative experiences. Sharing these experiences requires vulnerability, which can either be risky or beneficial (Barta and Andalibi, 2021). The opportunity to present oneself on social media opens up new avenues for identity development, including experimenting with self-presentation and expression (Weinstein and Davis, 2017). However, online self-disclosure often involves risks and harms (Barta and Andalibi, 2021), and adolescents, particularly those who are sensitive to rejection and self-presentation (Yau and Reich, 2019), justifiably feel more cautious about self-disclosure (Ryan and Ryan, 2019). Explicitly delivering effort investment in communication (Kelly et al., 2017) may be helpful, but interactions on current social media revolve around “Likes” which often fail to provide the necessary support to make users feel safe in those interactions (Prinstein, 2023).
Collectively, this line of work suggests a need for online spaces that support adolescent peer interactions in ways that are developmentally sensitive, i.e., providing support for authentic self-disclosure that leads to relationship development. Yet, research suggests that, currently, social media platforms instead prey upon the fears that adolescents experience, such as those stemming from their heightened sense of self-consciousness (Bukowski et al., 1998). For example, photo sharing, a prevalent form of communication on social media platforms, can exacerbate self-presentation concerns (Yau and Reich, 2019; Zhao et al., 2008). The perception of attractiveness is often central to adolescents’ self-concept and self-esteem, particularly among girls. These social media sites, with their visually oriented nature, the capability to carefully curate and manipulate visual images, and the existence of a peer audience that provides reinforcing feedback (Choukas-Bradley et al., 2020), often promote a never-ending cycle of upward social comparison. This comparison is made with meticulously selected and edited photos of peers, celebrities, and “influencers” (Fardouly et al., 2017), leading to unhealthy social comparisons and an amplified emphasis on self-presentation. Furthermore, social media platforms frequently place a higher emphasis on metrics such as time on site and return visits than on the quality of relationships. This approach encourages their users to display carefully curated content for approval and partake in continuous social comparisons.
2.4. Online Authenticity and Technology Design
While certain affordances (Evans et al., 2017; Treem and Leonardi, 2013; DeVito et al., 2017) of online communication—like asynchronicity, editability, visualness, and publicness—encourage manicured self-presentation, other affordances can promote authentic self-disclosure. Platforms and features such as Snapchat, Instagram Story, and BeReal use ephemeral content, which despite potential context loss (Cavalcanti et al., 2017), grants users greater control over their narrative and privacy (DeVito et al., 2017). This ephemerality can reduce self-consciousness, encourages casual and authentic exchanges (Bayer et al., 2016; Choi et al., 2020; Kreling et al., 2021; Xu et al., 2016; Chiu and Yuan, 2021), and enables deeper interactions, playful interactions, and private engagements (Phua et al., 2017; Choi et al., 2020; Chiu and Yuan, 2021; Xu et al., 2016).
Besides ephemerality, the perceived degree of audience control and the (smaller) size of the network also support authenticity. Users often set aside online spaces for closer relationships, fostering authentic interactions within intimate friend circles (Vaterlaus et al., 2016; Piwek and Joinson, 2016; Kreling et al., 2021; Chiu and Yuan, 2021; Taber and Whittaker, 2020; Huang and Vitak, 2022). Finsta users, for example, reconfigure Instagram to establish smaller, more intimate environments, enabling the sharing of unfiltered and emotionally expressive content (Xiao et al., 2020). While this active audience management can be demanding, it effectively promotes authenticity on social media (Xiao et al., 2020). Furthermore, other affordances such as pseudonymity that increase the perceived level of trust in a network provider and its users—commonly found on platforms like Tumblr—have been shown to contribute to online authenticity, particularly allowing for marginalized individuals seeking a space for authentic self-disclosure (DeVito et al., 2018; Davis, 2023; Barta and Andalibi, 2021; Schaadhardt et al., 2023).
Previous studies in Human-Computer Interaction (HCI) and Computer-Supported Cooperative Work (CSCW) have also sought to explore new ways of supporting relationship-building and authentic self-disclosure online. For instance, an analysis by Stepanova et al. (Stepanova et al., 2022) investigated the design of 50 artifacts aimed at promoting sincere feelings of connection in technologically driven systems. Their research identified nine diverse design strategies that enable genuine connection, with affective self-disclosure being the most pertinent to our study. Numerous systems employing this approach have relied on biofeedback sharing (e.g., signaling physiological activity such as heart rate and its associated emotion to a partner) in dyadic communication settings (Qin et al., 2021; Mentis et al., 2014; Liu et al., 2021, 2019; Gervais et al., 2016; Frey et al., 2018). Alternative strategies cater to broader connections beyond peer relationships, such as linking with the unity of humanity (Howell et al., 2019), engaging with strangers (Mitchell and Boer, 2017), or maintaining distant relationships (Robinson et al., 2020). In addition, a substantial body of research exists that examines how technology can facilitate self-disclosure and relationship enhancement through the use of situational conversation prompts. For instance, Zheng et al. (Zheng et al., 2021) developed PocketBot, a chatbot that nudges users to resume, enliven, and deepen conversations. However, research in this area has primarily focused on using conversational prompts to support relationships with strangers (Hu, 2022).
Finally, several studies have explored BeReal from authenticity and/or relationship angles (Vanhoffelen et al., 2023; Tirocchi, 2023; Bulchand-Gidumal, 2024; Snyder, 2024; Maddox, 2023). First, Tirocchi (Tirocchi, 2023) discusses how BeReal’s authenticity aligns with the values of users aged between 20 and 23, showing that young people’s media choices, including BeReal, reflect a desire for authenticity amidst a landscape of shifting social norms and expectations. Maddox (Maddox, 2023) highlights how BeReal diverges from performance-based authenticity by emphasizing the promotion of “authenticity-as-realness” through spontaneity. The authors note how the app’s panoptic aspect and self-monitoring are what enable users to put realness before performance, a culture not seen previously on other social media platforms. Taylor (Taylor, 2023) analyzed BeReal’s designs and its creation of a “live networked public,” suggesting that the platform’s unique affordances foster what is referred to as “sporadic” authenticity that is significant for its branding against other social media apps. On the other hand, Snyder (Snyder, 2024) critiqued BeReal’s claim to authenticity by examining its user experience and rules that enforce an ‘always-on’ mentality, arguing that this approach increases the need for external curation instead of promoting true self-expression. Finally, Reddy and Kumar (Reddy and Kumar, 2024) explored how BeReal’s design features shaped what constitutes “authenticity,” finding that BeReal’s affordances align more with extemporaneous interaction than comprehensive self-presentation. Despite extensive research on BeReal and online authenticity, there remains a limited understanding of how BeReal’s design aligns with adolescent developmental needs, particularly regarding self-presentation strategies and relationship-building on social media. Our study fills this gap by analyzing BeReal’s approach beyond affordances, examining how its design cultivates specific cultural norms and expectations of authenticity that resonate with teens’ heightened sensitivity to peer perceptions while also exploring how these factors contribute to both engagement and potential toxicity in self-presentation practices.
3. Background
Table with descriptions of BeReal’s key features as of December, 2022. Design Description Image Daily notifications at random times Once at any moment of each day, the app sends out a notification with the message, “Time to BeReal. 2 min left to capture a BeReal and see what your friends are up to!” Figure 1a Two-minute posting window Clicking on the daily notification initiates a two-minute countdown or an upward counting of how late the post is, shown above the camera Figure 1b & c Retake counter Users can view the number of times any photo, including a friend’s photo, was retaken before it was posted Figure 1d Ephemerality BeReal posts, comments, and reactions remain visible to other users until the subsequent notification is sent on the following day, indicating that another post is due, while the posters themselves retain continuous access. Front and back photoshoot All BeReal posts require that users capture images from both the front and rear camera almost at the same time Figure 1e & f No filter The in-the-moment pictures taken on BeReal cannot be modified, and no filters are available. Users are not able to choose past images from photo albums either. Photo feedback After the first photo is taken with the front(back) camera and while users wait for the second photo to be taken with the back(front) camera, the platform provides feedback on the initial image taken. The feedback exerts subtle pressure encouraging users to show their face and smile, for example, “Don’t move and smiiiile,” and “Wait, wait, wait, now smile.” Figure 1g Realmojis Realmojis enable users to respond to posts with a photo of their own instead of a standard “like” or “heart.” There are six options: thumbs up, happy face, surprised face, heart eyes, crying while laughing, and a lightning bolt. The first five, or “pre-made” (P18) Realmojis, utilize stored photos that users have taken at some point to represent each of these five emotions. The lightning bolt is an “Instant Realmoji” that requires a new photo each time. Figure 1h & i Reciprocal posting Users on BeReal can only view other people’s posts if they have contributed their content for the day Figure 1j Chronological feed Unlike most mainstream social media platforms, BeReal’s feeds are strictly chronological. Consequently, users see posts from all their BeReal friends, regardless of how frequently they interact or their relationship strength. Friend count hidden BeReal does not highlight the number of friends a user has. Users cannot see the friend count of others. Concerning their own friend count, they will see how many friends they have until they surpass 50 friends, at which point their friend count will be displayed as “50+.” Friction for adding new friends The platform explicitly promotes the intentional addition of friends with statements such as “On BeReal connect with your close friends” or “Add your real friends on BeReal” beneath the friend request button. Unlike other platforms, which make numerous friend suggestions, BeReal only syncs local phone contacts and does not allow user searches. Figure 1k
BeReal incorporates several design elements that funnel users towards creating a space for sharing unfiltered photos with a closer circle of friends. For instance, BeReal users receive daily notifications that encourage frequent, low-stakes posting. These notifications are sent at random times each day, as determined by the system, prompting users to post photos from both their front and back cameras within a two-minute window. Although users can postpone their response to the notification and post at a later time, such posts are tagged as “late.” Notifications about these “late” posts are sent to other users, and the delay in posting is highlighted above the post, e.g., “14 hr late.” The number of retakes is also visible to friends. While taking the photo, users are nudged to show their faces with prompts like “Your friends will definitely prefer to see your face!” and “Umm, anybody here?” Users can react to posts with “Realmojis,” emojis created from photos of their own facial expressions. Users are unable to view others’ posts until they have shared their own BeReal, and posts in feeds are displayed in chronological order. BeReal does not feature like counts or friend/follower counts. While users can see their own number of friends, any number exceeding 50 is simply displayed as “50+.” The platform also explicitly promotes adding “real friends” on BeReal. A detailed overview of these features, as of December 2022, when the study interviews were conducted, is presented in Table 1.
4. Methods
We conducted semi-structured interviews with 29 current BeReal users. All participants had used BeReal for at least one week, had at least one friend on the platform, and had posted at least three posts. Interviews were conducted over a period of two months, with 14 interviews taking place at the end of December 2022 and an additional 15 at the end of January 2023. The procedure was approved by our Institutional Review Board (IRB).
4.1. Participants and Recruitment
We used a combination of convenience and snowball sampling to recruit participants. Initial recruitment efforts involved reaching out to any known BeReal users aged between 13 and 18, as well as contacting the university’s research participation pool to seek potential participants. Three participants were recruited from this effort. Recruiters were offered an additional $10 incentive for recruiting friends who met the study’s inclusion criteria. Three additional participants were recruited this way. To reach a broader audience, we posted ads on Instagram and Facebook targeting 13-18-year-olds residing in the United States. A total of 195 individuals responded and completed a screener survey. From this pool, 46 respondents met our inclusion criteria. We scheduled and conducted interviews with 29 of these respondents on a first-come, first-served basis.
The mean age of the participants was 15.6 years (SD=1.54), with 22 (75.9%) identifying as girls, 6 (20.7%) as boys, and one (3.4%) as non-binary. The participants’ self-described racial/ethnic breakdown included 11 (37.9%) White/Caucasian, 9 (31%) Asian, 3 (10.3%) African American, 2 (6.9%) Hispanic, 1 (3.4%) Asian/White, 1 (3.4%) Hispanic/White, and 1 (3.4%) Asian/African. On average, participants reported having used BeReal for 5 months (mean=4.8 months, sd=1.7 months) at the time of the interview, excluding two novice users who had used BeReal for less than one month. To denote novice users with less than one month of experience with BeReal, add an “n” suffix to their participant name (e.g., “P10” as “P10n”). Approximately 58% of BeReal users in the U.S. identify as females (Curry, 2022). We do not have the racial or ethnic data for the overall population of BeReal users; our sample was predominantly White and Asian. The full list of participant information is available in Table 2.
Table with information on the participant ID, age, gender, ethnicity, time since BeReal install, and number of friends on BeReal PID. Age Gender Ethnicity Time since BeReal install No. of friends on BeReal P01 14 Boy White/caucasian 3 months 30 P02 13 Girl White/caucasian 2 months 22 P03 14 Girl White/caucasian 3 months 20 P04 17 Girl White/caucasian 5 months 25-30 P05 13 Girl African American 2 months 20-25 P06 16 Girl White/caucasian 3 months 50+ P07 17 Girl White/caucasian 4 months 50+ (150)** P08n 17 Non-binary White/caucasian 1 week 1 P09 16 Girl Asian/Asian American 5 months 50+ P10n 18 Girl African American 1 week 8 P11 17 Boy White/caucasian 6 months 50+ (70-80) P12 17 Girl African American 4 months 50+ (70) P13 14 Boy White/caucasian 4 months 20-30 P14 16 Girl Asian/Asian American 4 months 20 P15 15 Girl Asian/Asian American 4 months 30 P16 13 Girl Asian/White 5 months 3 P17 15 Girl White/caucasian 6 months 30 P18 15 Girl Asian/Asian American 6 months 15 P19 17 Boy Hispanic 7 months 20-40 P20 18 Boy White/caucasian 7 months 50 P21 17 Girl Asian/Asian American 5 months 30-50 P22 15 Boy Hispanic 4 months 10-15 P23 16 Girl Asian/Asian American 6 months 50+ P24 16 Girl Hispanic/White 8 months 50 P25 17 Girl Asian/Asian American 8 months 70 P26 15 Girl Hispanic 5 months 50+ (100+) P27 13 Girl Asian/Asian American 6 months 50+ P28 15 Girl Asian/African 5 months 15 P29 17 Girl Asian/Asian American 2 months 10 * Novice users are denoted with (n). ** Numbers in brackets (N) for 50+ friends is the participants’ estimate of the number of friends they have.
4.2. Procedure
One member of the research team conducted 29 semi-structured Zoom interviews. The interviews lasted between 30 and 45 minutes. Participants received an Amazon gift card worth US$20 as a thank-you for their time. The interview prompts probed 1) participants’ general experiences with using BeReal, 2) their engagement with other social media platforms, 3) their motivations for using BeReal, 4) how they engage with its features, and 5) their perceptions of online authenticity (see Appendix A for the complete interview protocol).
4.3. Data Analysis
We conducted a reflexive thematic analysis (Braun and Clarke, 2019) of the 29 interview transcripts. We chose this method of analysis as it is a structured yet flexible approach that is effective at generating a nuanced view of the data, aligns with a constructionist paradigm, and is not tied to a pre-existing theoretical framework, which is appropriate for this exploratory work. We began the analysis with open coding in three phases over the course of 2.5 weeks. In the first phase, the first author performed line-by-line coding of the transcripts. Codes during this phase were descriptive and close to the data, for instance, “weaponizing authenticity,” “tech empowerment,” and “casual, low-pressure.” In the second phase, all five members of the research team individually coded the same transcripts line-by-line using the Delve Tool (del, [n. d.]). The aim of this phase was to scrutinize the data as closely and comprehensively as possible. The third phase involved synthesizing the codes from the first two phases to construct higher-level themes. The five authors used the Miro Board (mir, [n. d.]) tool for affinity diagramming to visually group related codes into potential themes. The first author then synthesized these potential themes into a single set of proposed themes. Examples of such themes included “features and affordances that encourage or discourage inauthenticity,” “cultures and norms of authenticity on BeReal,” and “adolescents’ interactions with the features and culture of authenticity.”
Over the next two months, the research team met every 1-2 weeks to discuss, refine, and iterate on the codes and the constructed themes. At the end of these iterations, we had a finalized codebook with 43 refined codes, which were grouped into four overarching themes. After finalizing the codebook, four authors, excluding the first author, divided the 29 transcripts among themselves and coded different transcripts. The first author coded all of the transcripts. This process ensured that at least two members of the research team reviewed each transcript. Once all transcripts were coded, the authors who coded the same transcripts met to discuss and resolve any coding disagreements.
4.4. Ethical Considerations and Researchers’ Positionality
As academic researchers focusing on teen social media behaviors, particularly their use of the BeReal app, we recognize our positionalities as adults with limited firsthand experience of the current social media landscape as it is lived and navigated by teenagers. No one on the research team has been an active user of BeReal, and our engagement is limited to only two friends on the platform. This distance from the research subject(s), both in terms of generational gap and practical experience, fundamentally shapes our approach to this research.
From the outset, we positioned ourselves not as experts but as learners, genuinely curious about the nuanced ways in which teenagers engage with social media platforms like BeReal. We approached our participants with openness, explicitly communicating our lack of affiliation with BeReal and emphasizing that honest, authentic responses were invaluable to our understanding. By framing ourselves as outsiders looking to be educated, we hoped to mitigate any power dynamics inherent in the researcher-participant relationship and encourage a more forthright sharing of experiences.
The ethical considerations in conducting research with minors on sensitive topics such as social media behavior were paramount throughout our study. We obtained both written and verbal consent from all participants 111Consent from their parents was waived due to the practicality of recruiting., providing options for anonymity and comfort, such as turning off cameras during interviews. Our communications, especially when seeking clarifications via email, were crafted with care to be clear, respectful, and mindful of the participants’ privacy and well-being.
The perspectives that participants shared with us suggest that our approach succeeds in yielding rich, authentic perspectives. Teens shared critical reflections on the norms of BeReal, admissions of judgment towards others, expressions of self-consciousness, and deviations from expected engagement patterns.
5. Results
5.1. How BeReal’s Design Encourages Authentic Self-Presentation
“I feel like not showing your face …that’s one of the parts of what BeReal’s used for. They use their front-facing camera and the back-facing camera…[and have] little prompts that pop up saying ‘smile’ and whatnot. It just implies that you should show your face, and then next photo is showing what you’re doing right now…Same thing with the reaction emojis, that you react with your face, and if you’re just not showing your face on the platform, it just defeats the purpose of that front and back camera being used and the reaction emojis being used.” - P22
For the most part, participants did not join BeReal because they were seeking a platform that promotes authentic self-presentation. Instead, participants said they joined for social reasons, saying things like, “all my friends were on it” (P03) or “my friend told me to” (P09). Others were intrigued by the platform after coming across it on Instagram or TikTok. However, participants shared that once they became BeReal users, the platform’s features influenced their self-presentation and led them to share frequently and authentically (specifically, to post unfiltered, unstaged photos of themselves). Participants identified four key ways that BeReal’s design promotes authentic self-presentation: discouraging staged posts, relieving posting pressure, encouraging interactions with smaller social circles, and maintaining a simple, mission-focused app design. We described each of these in more detail below.
5.1.1. Designs that discourage staging.
BeReal has many features not present on other social media platforms that explicitly discourage users from staging their photos. First, the random timing of notifications deters staged posts. As P22 said, “I think it’s just the whole concept of the app, BeReal—it randomly going off. It’s supposed to catch you off guard.” They explained that the two-minute posting window encourages spontaneous posts, saying things like, “you don’t have time to get ready or to make yourself look good …nobody can stage something in two minutes” (P25). Similarly, P03 commented, “You can’t really fake it because it’s like a two-minute thing.” Separately, participants explained that the retake counter further discourages staging by creating subtle pressure to minimize retakes. They said things like, “People didn’t want other people to think that they were retaking all their photos and stuff” (P26). If a face is not detected in the frame, the platform will nudge the user to show themselves, with messages like “Your friends will definitely prefer to see your face!” and “Umm, anybody here?” As P13 shared, “When you take the photo and it just puts, like, the text up on the screen like, like ‘great smile’ or whatever, or like if you take an all-black photo it says like, ‘where are you?’ so …it gives you a good impression of [what BeReal is about] just based on those kind of little details.”
Participants explained that these features funnel them into very specific self-presentation choices. Some participants shared that they retake photos only when their “eyes are closed” or they “accidentally put [their] finger in the camera” (P11), not due to self-consciousness. They also avoided posting blank images, aiming to “always [have] something there” (P26), because blank posts “defeat the whole point” (P20). They said they try to post as soon as they receive a notification because waiting to post renders it “kind of pointless” (P04). They described the way BeReal encourages reciprocity and mutual sharing, saying, “You can’t see anyone else’s BeReal until you post yours for the day,” likening it to an “exchange” (P04). Given this context, posting a blank photo—withholding a piece of their lives—while being able to view others’ posts seemed like“cheating” (P13), “hypocritical” (P13), and “not fair” (P17) to several participants. Further, posting blank photos was seen as undermining or defeating the “point” (P15) or “purpose” of the app (P22). Participants said that when they see others post blank photos, their reactions range from mild disappointment to more pronounced disapproval. P13 said that a proliferation of blank photos would “ruin it” for them. Collectively, participants made clear that BeReal’s features communicate specific expectations for how they should engage with the platform. These designed expectations lead them to post unstaged photos and expect other users to do the same.
5.1.2. Designs that take pressure off posting.
Unlike most mainstream social media platforms, where sharing edited and curated content is the norm, BeReal’s designs allow users to share photos without the pressure to curate their posts by making sharing more low-stakes. For example, participants described the app-defined cadence of the daily posting by saying, “I think it does feel different just because I’m doing it so frequently, and because it has this connotation of, ‘okay, this doesn’t matter”’ (P24). Ephemerality further lowers the stakes for engaging with others. P16 shared: “I don’t feel as pressured to make it look super nice because it’s like it’ll just be away in 24 hours.” Moreover, the effort Realmojis demand and their ephemeral nature reduce the pressure to collect them. Unlike other platforms where “likes” signify social status, on BeReal, they do not carry much weight. As P15 put it, “There’s like no likes…I don’t think people usually really compare [the number of Realmojis they received].”
The effect of these features is that even compared to other casual online spaces like Snapchat, Instagram Close Friend Stories, and Finsta (Xiao et al., 2020), participants found BeReal to be more random, unlike other platforms where users put more effort into curating their content. P18 shared that they post photos on BeReal that they would not share on their Finsta because they are “just too boring.” P04 stated that Close Friends Stories were also considered less random than BeReal posts, as they are primarily used for sharing funny or noteworthy personal events. Because all users share whenever “the app tells [them] to do so” (P25), users feel comfortable under the shared vulnerability to post the “least exciting parts” (P27) of their day. P23 clarified the effect of the frequent, “low-fidelity” posting on BeReal stating, “[on other social media platforms] you have a purpose in posting these [posts]. BeReal, you’re not supposed to have a purpose in posting. You’re just posting just because of a random notification, throughout the whole day, would just tell you to post. There isn’t that intent behind it.” One participant made an analogy that BeReal “feels a lot like a Polaroid”, whereas Instagram “would be a picture hung on your walls—perfect and made to show off” (P29).
5.1.3. Designs that encourage audience curation.
Context collapse inevitably leads to curated self-presentation, and BeReal approached this phenomenon by encouraging users to curate their audience instead of their posts. Specifically, BeReal’s features encourage engagement with closer acquaintances and friends. This approach makes it easier for users to post casual, daily content without the fear of judgment, fostering an environment that promotes casual sharing rather than curated self-presentation. By enforcing reciprocity, the platform encourages users to limit their BeReal circles to those with whom they feel comfortable sharing their daily lives. As a result, users typically have fewer friends on BeReal than they do on other platforms. P04 explained that they add only the people they “talk to regularly or that someone that I would want to stay in touch with” because they wish to feel “comfortable with [friends] seeing parts of [their lives] that are personal.” Also, according to P10, scrolling through posts from less familiar friends on BeReal can become “overwhelming and just sort of boring.” Therefore, users tend to curate their friends, adding only those “that they care about seeing what they’re doing every single day” (P04). Moreover, by hiding friend counts, BeReal discourages popularity contests that can hinder authentic sharing. P04 explained: “On BeReal there’s definitely the least amount of pressure because nobody can see how many friends someone else has. Like, I don’t even know how many friends I have exactly…the number really doesn’t matter so you can have as few or as many as you want.” Overall, participants expressed that their friend lists on BeReal, in general, are “pretty much invite-only” (P22).
These features, which encourage audience curation, result in BeReal being described as“friendly” (P26),“cozy,” “home-like” (P16), and“communal” (P04). P18 noted,“Very few people will see my BeReal…it feels like there’s less pressure.” Participants also said that the BeReal experience made them“feel connected” (P28) to other users and appreciated the platform’s capacity to centralize close friends for updates that felt like“check-ins” (P14). Participants viewed the sense of a“judgment-free” (P16) environment that feels like a “safe, close space” (P29) fostered by a carefully curated audience as a defining characteristic that sets BeReal apart from other social media platforms.
5.1.4. Designs that focus on a single purpose.
As of December 2022, BeReal maintained a minimalist interface design, featuring only elements that serve its single purpose of sharing candid photos. This minimalism clarifies BeReal’s mission of authentic sharing with the users. For instance, participants viewed the act of retaking a photo as somewhat discordant with the platform’s mission and explained: “the point of BeReal is like you take one photo and you’re done” (P07).
P15 emphasized the platform’s focused mission, suggesting that expanding BeReal’s feature set could dilute its essence: “For BeReal, I feel like it’s the type of social media that’s a bit more hard to expand because expanding it, I feel like they would have to give up the whole point of BeReal. It’s hard to have both, where it’s a very authentic app, and have people creating content because they want others to view it.” Adding more features to BeReal would make it less distinct from other social media platforms. For example, P09 shared that “if you were to take multiple pictures, it would kind of become Instagram.” P23 echoed this sentiment, saying that “Instagram is complicated because there’s edits” and that BeReal is “supposed to be plain, simple, minimalistic.” P22 felt that Instagram’s additional features create “social hierarchy” but explained that on BeReal “there’s not a whole lot …to be judgmental of.”
The simplicity of BeReal not only makes its purpose clear to users but also sets the platform apart from features like TikTok Now, which, despite its striking resemblance to BeReal, is perceived differently due to its integration with TikTok’s broader functionality. For example, P25 expressed that “I feel like most people know TikToks as those short 15-second videos, whereas BeReal’s its own thing. And so to merge [Now into TikTok] just feels a little bit more awkward and then more clunky, and then it just doesn’t feel as real as BeReal.” Also, P24 shared that the audience base in the apps is different; “TikTok has more functions, and I guess that you can’t really curate the friends that you interact with on TikTok as much as you can with BeReal.” TikTok Now is a part of TikTok, which to the participants has “always been a general video sharing app” and “it’s the same concept as BeReal but …they don’t talk as much about like being real, that’s not really their focus the same way that it is with the BeReal app” (P08n).
5.2. The Perceived Benefits of BeReal’s Design
5.2.1. Validating an unfiltered life
Most participants expressed appreciation for BeReal’s mission to promote more authentic self-presentation. For example, P12 stated that it is “hard to like, eliminate that kind of idea of everyone and wanting to always appear like perfect and put together, but I think BeReal is doing a good job at trying to like begin and combat adding that type of thing.” P02 reflected on the impact of the authenticity norm on BeReal, noting a shift in how they view social comparisons: “There’s a lot of girls in my grade who I follow. And, like, at school, they present themselves as these confident …always feeling fabulous …and then seeing them in bed eating snacks, watching Netflix, and you’re like they’re a normal human being …they aren’t perfect all the time.” This brings out a sense of “normalcy” to unfiltered sharing as P28 succinctly put it, and offers a view of “their daily life rather than their glamorized events” (P25).
The contrast between the parts of themselves that people share on BeReal and Instagram was particularly striking to participants. For example, P24 observed that BeReal is “breaking down the walls that Instagram has put up since it was launched. You’re seeing people at their best, and you’re seeing people edited. BeReal is trying to get people to understand that that’s not real.” This aligns with P06’s assessment, who shared that “all I see of them on Instagram is like traveling” and that they find it “refreshing” that on BeReal, they can see their friends “staying at home” (P06). P24 further explained: “If Instagram is a room of people at their best moment with heavy editing, BeReal is trying to get people at their worst moment with no chance of editing. BeReal is anti-Instagram, and Instagram is just people trying to show off their best selves.”
Participants viewed BeReal’s emphasis on an unfiltered life as not just beneficial in the immediate sense but as indicative of a broader potential shift in social media culture. For example, P25 described BeReal as “trying to bridge that gap between social media and real-life” by “bring[ing] more intimacy into the social media world,” and they enjoyed being a part of “a collective effort to like destigmatize like having unfiltered moments.” Further, P08 expressed the platform’s mission as “hopeful and inspiring …it was created to help teach people that like you don’t need to stress all the time about social media and being cool and like the way that people perceive you. It’s okay to be your real self.” P24 expressed appreciation for BeReal’s new take on social media by acknowledging how BeReal devalues overthinking and excessive value given to an Instagram post, stating “There’s so many people that overthink it and give just an Instagram post too much value and platforms like BeReal are it doesn’t matter. It’s just an app …That lightheartedness is really valuable in terms of social media.”
5.2.2. Motivating positive changes
A few participants also said that using BeReal encouraged them to be the person they aspire to portray online. For example, P02 stated: “I’ll feel guilty if I am not, like, not getting out of bed or if like, my BeReal comes along and I’m still in bed or something, but most of the time it’s, like, it’s nice like when I’m doing something productive, and then my BeReal comes along, and I’m like, ‘oh cool I get to be doing something interesting while I take my BeReal.”’ Rather than feeling overwhelmed by this, P02 stated that they feel “motivated” and that they want to be productive when their BeReal comes along. P24 echoed this perspective by saying that looking back through their BeReal posts served as “a little mirror” to them.
Moreover, many expressed that, despite initial self-consciousness over seemingly mundane or unflattering posts, BeReal has empowered them to embrace and share their genuine selves. For example, P11 shared that they “kind of like to lean into the discomfort because it feels more honest.” P18 shared similar sentiments about overcoming self-consciousness, stating, “Well, I think about how it is important to be realistic on social media, and I think about how most people who see it wouldn’t really care that much …, so I try to just ignore it and press ‘post’ and be casual about it.” P02 echoed this sentiment, saying that “BeReal allows you to be authentic and kind of share what you’re going through. In an open way and like you’re supposed to be vulnerable, maybe, or you’re supposed to be your true self. So I think BeReal in some ways is healthy for people.”
Participants also observed that the norms and efforts toward authenticity on BeReal shielded them from the negative impacts of social comparison and approval anxiety often experienced on other platforms. P23 recognized the app’s focus on reducing the anxiety generated by metrics such as “likes,” “I don’t really get that many reactions either, but I don’t really care, so that’s the point of the app. I shouldn’t really care for what people are seeing or thinking about me when I’m posting it. That’s the point.” Furthermore, several participants noted that the changes they experienced on the app were self-directed, attributing this to the platform’s enjoyable and casual design, as highlighted by one participant who described it as “fun” (P06). For example, P04 shared, “I don’t think I’ve ever felt, like, forced …I feel like the whole idea is kind of, you’re not judging people, like, it’s not meant to be like that. It’s really just for fun, and you’re only on there if you want to be on there.”
5.2.3. Combatting toxicity with simplicity
BeReal’s straightforward design, with its focus on authentic photo sharing, significantly reduces toxicity levels compared to platforms like Snapchat, Instagram, and TikTok. The minimal use of comments on BeReal—often the breeding grounds for toxicity on other platforms—was seen by participants as a key factor contributing to a healthier user environment. P20 shared: “People can be pretty rude on other social platforms. But on BeReal, it’s not like Twitter where it’s like you can talk to anybody anytime. You have to find the person, you have to friend them. You have to engage on their posts to be able to contact them. It’s not as loose as maybe Snapchat or Instagram or Twitter.”
Many participants also appreciated that, unlike other social media platforms, BeReal’s minimalist design does not promote prolonged usage. As P13 shared: “when I’m looking through it, there’s never really anything negative …I don’t feel sucked into it …I don’t feel I never get on the app and I’m like, oh God I’m going to be on here for another hour or even ten minutes really. I just know it’s something I can open and close without like being like, oh one more post, one more video or whatever.” This view also suggests that BeReal’s simplified, chronological algorithm-based feed empowers users with greater control over how and how long they engage with the app’s content.
5.3. The Perceived Downsides of BeReal’s Design
Although BeReal has successfully nudged users toward an authenticity norm that mitigates many existing issues associated with mainstream social media platforms, we identified two shortcomings of its design: (1) it is boring and (2), at times, toxic.
5.3.1. Creating artificial pressure to post
Participants shared that they often feel pressured to be authentic, sometimes even against their will. As P25 shared, “I definitely feel like I am forced …And so I get curious, and then I end up posting anyways to see what’s going on.” The level of pressure that unfiltered selfies caused varied across participants, where some experienced “sometimes there are times where I don’t have makeup on or I don’t look my best and stuff because I was at the gym or something, and so it makes me feel a little stressed out,” for example. P19 also illustrated this issue, stating, “You just get so demotivated from feeling like you have to be real, where you’re just like, ‘Well, I don’t even want to touch the app at all.”’ Several features, like the lack of filters, retake count, “late post” label, public display of reactions, and the implicit pressure to keep up with posting on a daily basis, were identified as stress triggers.
A few participants felt that while the authenticity norm on BeReal carries significance, they felt the application of BeReal’s “rules” (P27) should put user comfort as the priority. As P29 explained: “I think some people are very obsessed with the idea of keeping things real, keeping things authentic…But to me, I think that it’s okay. I think it’s okay to be a little bit less authentic online. Only for reasons though, if you have a reason for it, I think that it’s okay.” They believed it acceptable to deviate from BeReal’s norms to avoid personal discomfort. P09 explained: “If I’m self-conscious about it, I won’t post about it, and I’ll just post about something else.” Similarly, P16 expressed that when they were not in the mood, they would simply point the camera “to my ceiling and then one is at the desk.” The need for positive self-presentation was a particular cause of discomfort. P25 reflected: “I still have the pressure of trying to make myself look cooler on BeReal. So if I’m just sitting alone, just chilling, watching TikToks, I usually don’t post BeReal until I’m out [of] my bed or doing something else at least.”
5.3.2. Inciting new forms of toxicity
Despite its intentions, BeReal’s simplicity and focus on everyday moments have inadvertently fostered new avenues for toxicity, with users sometimes judging each other’s content as “uninteresting” (P16), “underwhelming” (P19), and “boring” (P26), either expressing these judgments directly to their friends or confidentially with us during the study interviews. While participants were divided over whether the uninteresting nature of the posts stemmed from the posts themselves or the individuals creating them, some participants felt that the lack of engaging content was more directly associated with the users. For instance, P27 remarked, “Some posts show people just lying in bed. I think, ‘That doesn’t seem fun.’ If they do this every day, it appears they have no exciting point in their day.”
Anticipating or encountering such judgments led some participants to feel an amplified fear of negative judgment when posting on BeReal, which usually involved more vulnerability than posts on other social media platforms in general. As P05 noted, “People could make fun of your true self.” This led some participants to “just walk outside and take a picture.” (P29). The participant further explained, “Even if they don’t necessarily go out, they wait and just leave the house a bit, open the door, take a picture, and then just walk back inside.” P20 further pointed out the sometimes paradoxical nature of the authenticity norm on BeReal: “People might judge you negatively because they compare you to the people who aren’t authentic.” Since it is possible that not everyone embraces BeReal’s “authenticity” to the same extent, being authentic could sometimes lead to negative consequences. P25 shared a candid example: “I know for me, we all glamorize those super social party raves type of people who hang out, …Whereas I feel like when I try to be authentic, it’s like I don’t do that and usually it’s just me at home watching my shows. And so, it definitely conflicts with what society expects versus what I want to do with my life.”
Another particularly salient issue arises when users begin to critique each other’s level of authenticity. P17, for instance, overheard a friend remarking, “Oh, look, they just retook their BeReal 12 times,” while P12 shared a similar story about a conversation her sister had over the phone: “Oh my gosh this person retook their BeReal like several this many times, and I only did it this many times can you believe it?” This observation was not isolated, and participants P02, P06, P11, and P20 noted similar interactions.
5.3.3. An underwhelming user experience
BeReal’s focus on authenticity ironically results in a dull and monotonous user experience. As P19 noted, the platform’s mandate of authenticity, combined with specific posting times and limited features, feels “overbearing.” Ironically, the platform’s stringent structure and time constraints contribute to repetitiveness and ennui, rendering the overall user experience “underwhelming” (P19) for some participants. Participants also shared that the scarcity of compelling features amplifies the platform’s tedium. BeReal has been likened to a fleeting craze, with users reflecting, “It’s one of those fads, in a sense” (P19), and observing, “I think… BeReal kind of died, and it was just a fad” (P26). It appears that BeReal’s intensely authenticity-driven design cultivates an “underwhelming” user experience, pushing users away from the platform even as it strives to pull them in with its authentic appeal.
5.4. BeReal’s Potential and Limitations to Support Relationship-Building
Participants also reflected on the extent to which BeReal’s encouragement to share unstaged photos at unexpected times might enable relationship-building (or fail to do so). We examined these reflections from the perspective of Social Penetration Theory and reported on teens’ thoughts about BeReal’s potential for increasing the breadth and depth of self-disclosure.
5.4.1. Enabling casual relationship building
Participants noted how BeReal provides insights into the daily lives of friends, offering a more intimate view of moments they might not otherwise share. P21 gave an example: “I can see things that I normally wouldn’t have known about [my friends]. My friend the other day, I didn’t know this, but she got boba at the most random time …it gives me a new insight into what people do in their lives when I’m not with them, so that’s why I can appreciate it.” In this way, BeReal can support relationship building by broadening users’ perspectives about others.
Many of our participants also expressed that BeReal is a valuable tool for maintaining connections and staying in touch with their friends. They specifically valued the ability to catch a glimpse of the day-to-day activities of friends who live far away. P04 shared: “I get to see, like, a little bit of my, some of my friends’ days that I wouldn’t see otherwise because they don’t live in [city name] they live in other places …I wouldn’t normally see a picture of them at school or something like that, but it’s kind of, it’s kind of nice to see …a little glimpse into people’s lives.”
5.4.2. Limited potential for deepening relationships
Many of our participants reflected that they want to share different facets of themselves online, suggesting a lack of nuance in the kinds of self-disclosure that BeReal makes possible. They found it limiting that on BeReal they are not able to share “[their] energetic style or [their] talkative style” (P28), that they “like listening to K-pop” (P09), “a screenshot…of a game [they’re] playing or a song [they’re] listening to” (P16), and “thoughts of the day or whatever you’re doing” (P14), for example. They wanted to be able to share “anything from like my camera roll that I took like maybe like ten minutes ago” (P03), or just “a lot of things” (P15) that they would share or see on their Instagram, Finsta in particular, online.
Our participants also shared that what they post on BeReal is “very surface level” (P15) since it is “just a photo” (P05), implying that depth of self-disclosure is also lacking. They expressed that they are not able to share their emotions (P02, P16), personality (P18), values (P27), and political beliefs (P06, P15) as they do on other platforms such as Instagram (P06, P09, P18), Snapchat (P09), or Twitter (P02). They also confirmed that the lack of such varied forms of exchange makes it harder for them to learn more about and feel closer to other users on BeReal than on other platforms where deeper self-disclosures happen. P16, for example, shared: “I think you can’t really communicate your emotions that easily through [BeReal], which to me is a pretty important part of getting to know somebody, what kind of things they’re going through.” Such limitations led participants to suggest BeReal adopt other forms of content-sharing, including live photos or text posts.
5.4.3. An inability to share with specific others
Though BeReal encourages users to curate their audience, it was difficult for some of our participants to draw that boundary given their offline social dynamics. As Social Penetration Theory explains, people build relationships by disclosing intentionally within the context of a specific relationship. Users’ inability to target who they shared with exacerbated their feeling of intrusion or overwhelmed with making BeReal posts. For example, P17 shared, “I’m in high school, and so a lot of it is just people in my grade that I wouldn’t necessarily trust. …So there have been times where I’ve retaken my BeReal, because I didn’t like the way I looked in a photo. Which I think is not the point or what they originally intended, but I think that in the social situation, like high school is, I feel like that it was bound to happen in a way. And I know people that’ll retake their BeReals 10 times just to get the right angle.” They further emphasized the limits of sharing “authentic” content if they are not close with their BeReal friends, stating, “In my opinion, I think no one is really going to see the authentic you but yourself, or maybe someone you’re really close with, like a partner or a really close friend or really close family member. Because everyone around you isn’t going to be able to see you truly unless they know you really well.”
Similarly, some participants perceived unfiltered photo-sharing to be relatively high-stakes and intimate. For example, P25, who called themself a “homebody”, expressed that they do not want other people to see that they “stay at home all the time.” Such feelings of intrusion were especially evident when participants had friends on BeReal with whom they were not very close. For example, P18 shared, “I do also worry about what I look like, because there’s people that I’m not that close with, that I’m friends with on there, so I care about what I look like, and I cut off my face if I feel like I look bummy.”
5.4.4. Coerced self-disclosure cannot enable relationship building
Some participants explained that for self-disclosure to be authentic, it must be self-motivated. P29 expressed, “the most important thing is that someone has to be comfortable with [sharing]” and that “if you’re uncomfortable with something, then you’re not being real, per se,” suggesting self-motivation as a prerequisite for authenticity. P19 also shared that BeReal is “technically authentic, but at a price.” They elaborated, “I guess I could call it real, but at the same time, I feel like there’s much more just that isn’t worth it. You get self-conscious, feel the need to do it, you feel like you have to do it.”
Moreover, our participants felt that intentions should underpin the authenticity of self-disclosure. As P19 exemplified, even staged social media photos can be “perfectly authentic” if the user consciously acknowledges their staging. Conversely, an ‘in-the-moment’ BeReal with no staging can be considered inauthentic if the disclosure lacks self-motivation or intentionality. Hence, even in the face of modern social media norms where staging equates to inauthenticity, intentional staging can maintain authenticity in a user-defined sense, even if not communicated to the audience. On a related note, personal significance was also often mentioned as a factor not reflected in BeReal’s platform-defined authenticity norms. P29, for example, reflected: “If you are holding a post and you choose to post later…I think to me is authentic. That shows that this was something important to you…I think it has to mean something to you first, to be authentic.” Conversely, BeReal posts shared without curation were perceived as lacking the depth of self-disclosure necessary for meaningful social connection.
5.5. Adolescents’ Conclusions about BeReal
5.5.1. Bound to fail
Some participants expressed cynicism about the potential for BeReal to achieve its mission of encouraging authentic self-presentation, explaining that the platform cannot escape the fact that on any social media platform, users will seek approval and present overly positive versions of themselves. For example, P17 shared, “Yes, people are showing themselves kind of unedited, but I think also it was kind of bound to fail in some ways. No matter what, people are going to want to impress other people. No matter what, people are going to want to have other people see them look good.” They added “I think any social media platform, no matter what it is, it’s never going to be the authentic …because …if you’re looking for someone else’s approval, that’s not being authentic,” suggesting that authenticity and social media are fundamentally paradoxical. P25 shared a related perspective that online self-disclosure should be approached with caution, stating, “yeah, it is promoting authenticity, but you still have to remind yourself that this is still online, and so your digital footprint is still there, and everything you post will still be there, and people will see that. So don’t get too sucked in. Don’t glamorize it too much because it’s still something that’s on the phone and something. It’s still considered a social media.”
Some of our participants also asserted that motivations to preserve private experiences precede self-disclosure on social media platforms. P25 captured this sentiment, saying that displaying everything on social media, including mundane activities like “enjoying a cup of fruit by yourself”, can cause users to “lose a lot of the special aspects” of their lives. They shared additional insight: “I feel like if we’re too authentic on social media, it takes away the beauty of being connected in real life. If everybody knows what you’re doing all the time, then there’s no real privacy or connection between the people who see you every day because your every day is already on there too.” P20 also pointed out that certain aspects of life should be kept private, stating,“you’re not going to post your entire life on social media unless you’re a vlogger or an influencer. But even then, you still have parts where you don’t want to show other people.”
5.5.2. Overly prescriptive
Some participants outright rejected the idea that they should adhere to BeReal’s nudges, viewing them as either unnecessary pressure or misaligned with their social media usage. P14 epitomized this perspective, stating that BeReal serves primarily as “pure entertainment” and that people use the platform as an “excuse to take a photo and share it.” P25 went a step further, expressing clear discomfort with an overly serious approach to the app’s rules, stating, “If someone…need[s] to do a couple retakes within the two minutes, who gives a crap? Even if it’s not real, but I think everybody’s just taking it too seriously.”
Participants reported that they have plenty of options for flouting the platform’s demands. For example, users can choose to wait before taking photos, and users can technically retake photos infinitely. This flexibility meant that the unconvinced users did not follow BeReal’s norm at all, providing them justification for staging or delaying posts for curated self-presentation. For instance, P09 shared, “Personally, I kind of ignore the fact that they tell me to be more real and I kind of just take a photo whenever I feel like.” P05 echoed this view: “I feel the same as like Instagram because, like, you could honestly post whenever you wanted.”
5.5.3. Nevertheless, a welcomed change
Despite criticisms, the consensus among participants is that BeReal’s initiative to foster authenticity is “a great thing” (P22)—a positive departure from the norm of other social media platforms. For example, P06 stated that they appreciated that BeReal “wants to give authenticity” to its users because “it’s never been done before and so it’s something different to see and …experience.” P29 also shared that they like the “ overall message of what [BeReal is] trying to do” as they feel that BeReal “encourages people to not fall into social media’s expectations for what they should and shouldn’t do, or how their life should and shouldn’t be.” P02 went further, suggesting that “it would be amazing if those apps [i.e., Instagram], or maybe popular people on those apps, would start trends where you were more authentic. So when you got your BeReal, your BeReal notification you would also post something on Instagram to show like vulnerability on all social media apps.” For those who felt dissatisfied with BeReal’s take on “authenticity,” they felt that BeReal is “less inauthentic” (P10n) at the very least, and that it is “the most authentic photo version of yourself” given BeReal photos indeed cannot be edited.
6. Discussion
Peer relationships play a crucial role in adolescent development, and reciprocal and authentic self-disclosure is necessary for these relationships to develop. Despite the significant role of social media in facilitating such adolescent peer interactions, its capacity to support authentic self-disclosure is often compromised, for example, by context collapse and the pressures of self-presentation. This study delves into BeReal, an app distinguished by its focus on fostering reciprocal, authentic self-presentation.
Our research examines BeReal’s design in supporting authentic self-presentation on social media and assesses its alignment with the developmental and self-presentation needs of teens. Here, we (1) introduce the Social Media Self-Presentation Matrix (SMSM) based on literature to guide the specific design goals in supporting self-presentation online, cognizant of the different goals and processes of self-presentation, (2) discuss the implications of social media design based on the matrix, and (3) propose design recommendations for authentic sharing on social media derived from our study findings.
6.1. The Social Media Self-Presentation Matrix (SMSM)
A 2x2 matrix with two types of self-presentation goals on the horizontal axis: “Self-Enhancing” (i.e., strategic social maneuvering goal) and “Self-Expressive” (i.e., accurate self-portrayal goal). On the vertical axis, we categorize the cognitive process of self-presentation as either “Automatic” (i.e., without conscious awareness, with minimal cognitive effort) or “Controlled” (i.e., deliberate, planned). The quadrant of Controlled and Self-Enhancing self-presentation is exemplified by the general strategy of highly manicured self-presentation on broadcasted social media platforms like Instagram. BeReal is intended to fall under the Automatic, Self-expressive quadrant. We identify a design opportunity for social media platforms to move toward the Controlled, Self-expressive quadrant. This quadrant would represent a more realistic approach toward authenticity, considering the intricacies of social media interactions, such as privacy issues, and the developmental characteristics of teens, such as heightened self-consciousness. We anticipate that a shift toward this quadrant could, in the long run, lead users or platform norms toward Automatic, Self-expressive self-presentation, which is what BeReal aims to achieve.
6.1.1. Evaluating BeReal’s Design Support for Self-Presentation Using the SMSM
Drawing on theories of self-presentation (Schlenker, 2003), we identify two axes for describing self-presentation on social media. The first axis, Automatic vs. Controlled, captures the cognitive processes that drive self-presentation. The second axis, Self-Enhancing vs. Self-Expressive, captures the goals of self-presentation. Combining these axes, we create a matrix that describes four modes of self-presentation strategies potentially dictating social media platform designs (see Figure 2). We use this matrix to evaluate BeReal’s design in supporting authentic self-presentation and suggest alternative ways for achieving BeReal’s design goals.
Social media platforms generally promote a Controlled, Self-Enhancing mode of self-presentation, but BeReal has attempted to shift towards a more Automatic, Self-expressive model. This effort has been partially successful due to several design choices, such as but not limited to: randomly timed notifications that discourage staged sharing, daily ephemeral posting that reduces posting pressure, hidden follower counts that support audience curation, and a minimalist design that underscores the platform’s single-minded focus on extemporaneous, unfiltered photo-sharing. These elements collectively fostered user behaviors distinctly different from those on platforms like Instagram and were largely welcomed by users.
However, challenges remain. For example, users have reported feeling new kinds of pressure and discomfort, leading to doubts about the feasibility of authentic self-presentation in social media contexts. Despite BeReal’s efforts, some users felt the platform did not support their relationship-building needs, finding it difficult to present themselves accurately through the constrained format of an uncurated, unfiltered photo.
6.1.2. Challenges and Considerations for Achieving Automatic, Self-Expressive Self-Presentation on Social Media
A closer examination of self-presentation theory, as discussed in Section 2, may clarify the mismatch between BeReal’s ideal and the reality experienced by teens. BeReal aimed to transition users’ general self-presentation strategies from a Controlled, Self-Enhancing model to an automatic, self-expressive one. Achieving this shift requires several conditions, such as a familiar and supportive audience group, minimal discrepancy between one’s self-concept and presented self, adequate skills and resources for accurate self-presentation, and low stakes in general, which would then allow for a more Automatic approach.
However, several factors made this shift challenging. Some users had large friend groups that included people they did not fully trust, making Automatic self-presentation unrealistic. Others felt more self-conscious because the platform nudges them to reveal aspects of their lives they typically choose to conceal, such as social reticence. Further, BeReal restricts users to a single, unfiltered photo at a randomly assigned time with a two-minute window. This demands significant skill for accurate representation relative to self-presentation on other platforms with no such constraints. Furthermore, the general negativity in the perception of social media exemplified by privacy concerns tends to guide self-presentation behavior in a more Controlled direction rather than an Automatic one.
Most importantly, self-presentation on social media is inherently guided by external validation and cues, demanding a high level of maturity to balance personal and inner concerns with the outer ones effectively—a level of sophistication that may be beyond what can typically be expected of teenagers (Schlenker, 2003). Indeed, we observed instances where BeReal’s design features inadvertently prompted adolescents to overlook their internal cues about whether to disclose or withhold, at times pressuring them to share when they would prefer not to.
These observations suggest that transitioning from Controlled, Self-Enhancing self-presentation to Automatic, Self-Expressive self-presentation is not straightforward due to inherent challenges. Social media inherently involves broader social interactions than F2F interactions, meaning that even with a smaller friend group, the audience is larger, making Automatic self-presentation difficult. Additionally, the prevalence of self-enhancement on social media platforms in general makes Self-Enhancing self-presentation the default strategy. However, Automatic self-presentation processes typically emerge in routine situations that are frequently encountered. Therefore, encouraging Self-Expressive self-presentation on social media requires creating opportunities for more Controlled processes to coexist, allowing users to navigate the complex dynamics of online social interactions more deliberately and effectively.
6.2. Design Recommendations to Support Authentic Self-Presentation Online
Informed by data derived from our study and SMSM, we propose design recommendations for fostering authentic self-presentation on social media.
Support “accurate” self-presentation. As seen in Section 2, self-presentation can be driven by truthful motives, such as the desire to ensure others see them as they truly are. However, accurate self-presentation often requires skill and deliberation. Therefore, providing users with the time and opportunities for this is a crucial aspect of authentic self-presentation on social media. For example, platforms could offer daily prompts or missions that encourage authentic reflection and sharing, allowing users to explore the breadth and depth of self-disclosure. Providing more varied modes of expression can additionally facilitate the Automatic aspects of self-presentation, particularly when it enables self-presentation that aligns with consistent self-images and personality traits.
Provide air cover for audience management. Creating a safe space is essential for facilitating automatic self-presentation. While Automatic processes may seem less relevant to authenticity on social media, it is important to recognize that Controlled self-presentations require significant effort and can be mentally exhausting, potentially leading to self-regulation challenges. This exhaustion not only increases problematic social media use but may also diminish users’ effectiveness in managing their subsequent self-presentations, such as oversharing or appearing arrogant. By protecting the cognitive resources of users, particularly children, we can enhance their capacity for more accurate and authentic self-presentation.
Scaffold reciprocal sharing. According to SPT, mutual disclosure is an essential component of social penetration progression. However, lurking behaviors are prevalent on many mainstream social media platforms. To normalize reciprocity and foster deeper social interactions at an individual or community level, designers might consider introducing mechanisms to minimize the vulnerability associated with initiating online engagement. As reported by our participants, daily post notifications acted as a support mechanism for posting and relieved the pressure associated with voluntary sharing on social media. Thus, strategies such as daily reflection prompts might also serve to diminish the vulnerability associated with initiating disclosures and encourage more users to reciprocate self-disclosure.
Prevent weaponized authenticity. Participants said that authenticity norms turned toxic when the platform pressured them to over-disclose or enable users to shame one another, police one another, or compete to be the most authentic. Through intentional design decisions, platforms can encourage candid sharing without making an idol of it. For example, a platform might avoid quantifying authenticity by choosing not to display the number of times a photo was retaken.
It is important to note that while self-disclosure is generally desirable for social penetration, pursuing it online could introduce potential risks or unforeseen consequences not typically present in offline settings. Younger users, especially those with problematic smartphone use (Davis et al., 2023) are even more likely to disclose beyond what may deemed the appropriate level of self-disclosure. Therefore, when designing to expand the breadth, depth, or frequency of online self-disclosure, it is essential to anticipate potential harmful outcomes to users, such as online harassment or impersonation. Designers should rigorously test their designs with a diverse user group to ensure no targeted group faces heightened risk, implement robust privacy and security measures on the platforms (e.g., disabling screenshots), and possibly guide users to understand and reflect on their desired level of self-disclosure and/or social penetration online.
6.3. Limitations and Future Work
Our study has several limitations. The first is that our participant sample includes significantly more people identifying as girls than as boys or non-binary. Further, BeReal introduced many new features since our data collection: (1) the integration of Spotify, where the cover art of the song or podcast a BeReal user is listening to will appear in the bottom right corner of their post, (2) a bonus BeReal where BeReal allows users to post an additional two BeReals throughout the day — but only if they post their first during the allotted two-minute window, (3) branded accounts (e.g., Adidas, Notion), and (4) “Groups” which resemble group chats, to name a few. A lot of how the platform norm is perceived is likely to have changed with the introduction of the new features. It would be useful for future work to explore how this extra disclosure and extra element of gamification support or undermine healthy and desired forms of authentic self-presentation or create further privacy concerns or pressure.
Future studies might also consider exploring methods to support more comprehensive authentic sharing integral to adolescent relationship-building. Specifically, while BeReal’s design indeed offers a new approach to social media—encouraging users to share more authentically—our study’s participants highlighted that the platform focuses primarily on one facet of self-disclosure: transparent photo sharing. They also mentioned the presence of potentially harmful, undesired, or incomplete forms of authenticity on BeReal. As a result, there remains a need for future research to explore novel designs that could cultivate different modes of sharing that transcend a single form of self-disclosure or self-presentation on social media, moving towards, for example, more intimate self-disclosure or deliberate self-presentation that our participants expressed a desire for.
7. Conclusion
In this paper, we explore the impact of design on promoting authenticity in social media. We conducted a qualitative interview study with 29 participants who use BeReal, a social media platform created to reduce self-presentation pressures and inauthenticity that arise from excessive content curation and strategic self-presentation. Our findings demonstrate that the affordances of BeReal effectively generate a space for authentic self-presentation, where users share their unfiltered lives and engage in relatively little social comparison. However, we also find that transparent photo sharing is only a single facet of our teen participants’ understanding of authentic self-presentation, and young people are eager for online spaces that allow them to share more than just the details of their day. This study contributes to the important discourse about the influence of social media on teen relationship-building and presents a framework and guideline for developing platforms that support the different relationship and self-presentation needs of the youth.
Acknowledgements.
The authors would like to acknowledge the CERES network, which provided support for this work. We additionally thank the anonymous reviewers for their detailed feedback and the participants for sharing their thoughts. We truly appreciate all the help.References
- (1)
- noa ([n. d.]) [n. d.]. BeReal. Your friends for real. https://apps.apple.com/us/app/bereal-your-friends-for-real/id1459645446. Accessed: 2023-7-11.
- del ([n. d.]) [n. d.]. Qualitative Data Analysis Software. https://delvetool.com/. Accessed: 2023-7-15.
- mir ([n. d.]) [n. d.]. The Visual Collaboration Platform for Every Team. https://miro.com/. Accessed: 2023-7-18.
- noa (2023) 2023. Social Media and Youth Mental Health: The U.S. Surgeon General’s Advisory. Technical Report.
- Aacap ([n. d.]) Aacap. [n. d.]. Social Media and Teens. https://www.aacap.org/AACAP/Families_and_Youth/Facts_for_Families/FFF-Guide/Social-Media-and-Teens-100.aspx. Accessed: 2023-7-17.
- Anderson et al. (2023) Monica Anderson, Michelle Faverio, and Jeffrey Gottfried. 2023. Teens, social media and technology 2023. Pew Res Cent (2023).
- Argyle (2013) Michael Argyle. 2013. The psychology of happiness. Routledge.
- Bailey et al. (2020) Erica R Bailey, Sandra C Matz, Wu Youyou, and Sheena S Iyengar. 2020. Authentic self-expression on social media is associated with greater subjective well-being. Nat. Commun. 11, 1 (Oct. 2020), 4889.
- Bargh (1996) John A Bargh. 1996. Automaticity in social psychology. (1996).
- Barta and Andalibi (2021) Kristen Barta and Nazanin Andalibi. 2021. Constructing Authenticity on TikTok: Social Norms and Social Support on the “Fun” Platform. Proc. ACM Hum.-Comput. Interact. 5, Cscw2 (Oct. 2021), 1–29.
- Baumeister (1982) Roy F. Baumeister. 1982. A self-presentational view of social phenomena. Psychological Bulletin 91 (1982), 3–26. https://doi.org/10.1037/0033-2909.91.1.3 Place: US Publisher: American Psychological Association.
- Bayer et al. (2016) Joseph B Bayer, Nicole B Ellison, Sarita Y Schoenebeck, and Emily B Falk. 2016. Sharing the small moments: ephemeral social interaction on Snapchat. Inf. Commun. Soc. 19, 7 (July 2016), 956–977.
- Blumer (1986) Herbert Blumer. 1986. Symbolic interactionism: Perspective and method. Univ of California Press.
- Boyd (2008a) Danah Boyd. 2008a. Why Youth (Heart) Social Network Sites: The Role of Networked Publics in Teenage Social Life. In The John D. and Catherine T. MacArthur Foundation Series on Digital Media and Learning. unknown, 1–26.
- Boyd (2008b) Danah Michele Boyd. 2008b. Taken out of context: American teen sociality in networked publics.
- Braun and Clarke (2019) Virginia Braun and Victoria Clarke. 2019. Reflecting on reflexive thematic analysis. Qualitative research in sport, exercise and health 11, 4 (2019), 589–597.
- Brown (2014) Jonathon Brown. 2014. The self. Psychology Press.
- Bukowski et al. (1998) William M Bukowski, Andrew F Newcomb, and Willard W Hartup. 1998. The company they keep: Friendships in childhood and adolescence. Cambridge University Press.
- Bulchand-Gidumal (2024) Jacques Bulchand-Gidumal. 2024. The case of BeReal and spontaneous online social networks and their impact on tourism: research agenda. Curr. Issues Tourism 27, 1 (Jan. 2024), 7–11.
- Cacioppo and Cacioppo (2014) John T Cacioppo and Stephanie Cacioppo. 2014. Social relationships and health: The toxic effects of perceived social isolation. Social and personality psychology compass 8, 2 (2014), 58–72.
- Carpenter and Greene (2015a) Amanda Carpenter and Kathryn Greene. 2015a. Social Penetration Theory. In The International Encyclopedia of Interpersonal Communication (1 ed.), Charles R Berger, Michael E Roloff, Steve R Wilson, James Price Dillard, John Caughlin, and Denise Solomon (Eds.). Wiley, 1–4. https://doi.org/10.1002/9781118540190.wbeic160
- Carpenter and Greene (2015b) Amanda Carpenter and Kathryn Greene. 2015b. Social penetration theory. The international encyclopedia of interpersonal communication (2015), 1–4.
- Cavalcanti et al. (2017) Luiz Henrique C B Cavalcanti, Alita Pinto, Jed R Brubaker, and Lynn S Dombrowski. 2017. Media, Meaning, and Context Loss in Ephemeral Communication Platforms: A Qualitative Investigation on Snapchat. In Proceedings of the 2017 ACM Conference on Computer Supported Cooperative Work and Social Computing. Acm, Portland Oregon USA, 1934–1945.
- Cheek and Hogan (1983) Jonathan M Cheek and Robert Hogan. 1983. Self-concepts, self-presentations, and moral judgments. Psychological perspectives on the self 2 (1983), 249–273.
- Chiu and Yuan (2021) Hsuen Chi Chiu and Chien Wen (tina) Yuan. 2021. To Last Long or to Fade Away: Investigating Users’ Instagram Post and Story Practices. In Companion Publication of the 2021 Conference on Computer Supported Cooperative Work and Social Computing. Acm, Virtual Event USA, 32–35.
- Choi et al. (2020) Sukyoung Choi, Dmitri Williams, and Hyeok Kim. 2020. A snap of your true self: How self-presentation and temporal affordance influence self-concept on social media. New Media & Society (Nov. 2020), 146144482097719.
- Choukas-Bradley et al. (2020) Sophia Choukas-Bradley, Jacqueline Nesi, Laura Widman, and Brian M Galla. 2020. The Appearance-Related Social Media Consciousness Scale: Development and validation with adolescents. Body Image 33 (June 2020), 164–174.
- Cohen (2004) Sheldon Cohen. 2004. Social relationships and health. American psychologist 59, 8 (2004), 676.
- Curry (2022) David Curry. 2022. BeReal Revenue and Usage Statistics (2023). https://www.businessofapps.com/data/bereal-statistics/. Accessed: 2023-7-12.
- cycles and Text ([n. d.]) This text provides general information Statista assumes no liability for the information given being complete or correct Due to varying update cycles and Statistics Can Display More up-to-Date Data Than Referenced in the Text. [n. d.]. Topic: BeReal. https://www.statista.com/topics/10096/bereal/
- Darr and Doss (2022) Christopher R Darr and Erin F Doss. 2022. The Fake One is the Real One: Finstas, Authenticity, and Context Collapse in Teen Friend Groups. J. Comput. Mediat. Commun. 27, 4 (July 2022), zmac009.
- Davis (2012) Katie Davis. 2012. Friendship 2.0: adolescents’ experiences of belonging and self-disclosure online. J. Adolesc. 35, 6 (Dec. 2012), 1527–1536.
- Davis (2023) Katie Davis. 2023. Technology’s child: digital media’s role in the ages and stages of growing up. The MIT Press, Cambridge, Massachusetts.
- Davis et al. (2023) Katie Davis, Petr Slovak, Rotem Landesman, Caroline Pitt, Abdullatif Ghajar, Jessica Lee Schleider, Saba Kawas, Andrea Guadalupe Perez Portillo, and Nicole S Kuhn. 2023. Supporting Teens’ Intentional Social Media Use Through Interaction Design: An exploratory proof-of-concept study. In Proceedings of the 22nd Annual ACM Interaction Design and Children Conference (Chicago, IL, USA) (Idc ’23). Association for Computing Machinery, New York, NY, USA, 322–334.
- DeVito et al. (2017) Michael A DeVito, Jeremy Birnholtz, and Jeffery T Hancock. 2017. Platforms, People, and Perception: Using Affordances to Understand Self-Presentation on Social Media. In Proceedings of the 2017 ACM Conference on Computer Supported Cooperative Work and Social Computing. Acm, Portland Oregon USA, 740–754.
- DeVito et al. (2018) Michael A DeVito, Ashley Marie Walker, and Jeremy Birnholtz. 2018. ’Too Gay for Facebook’: Presenting LGBTQ+ Identity Throughout the Personal Social Media Ecosystem. Proc. ACM Hum.-Comput. Interact. 2, Cscw (Nov. 2018), 1–23.
- Elkind (1967) D Elkind. 1967. Egocentrism in adolescence. Child Dev. 38, 4 (Dec. 1967), 1025–1034.
- Erikson (1994a) Erik H. Erikson. 1994a. Identity: Youth and Crisis. W. W. Norton & Company. Google-Books-ID: nGqc6JxV0aQC.
- Erikson (1994b) Erik H Erikson. 1994b. Identity: Youth and Crisis. W. W. Norton & Company.
- Evans et al. (2017) Sandra K Evans, Katy E Pearce, Jessica Vitak, and Jeffrey W Treem. 2017. Explicating Affordances: A Conceptual Framework for Understanding Affordances in Communication Research: EXPLICATING AFFORDANCES. J Comput-Mediat Comm 22, 1 (2017), 35–52.
- Falk et al. (2014) Ph.D Falk, Emily B., M.A Cascio, Christopher N., Ph.D Brook O’Donnell, Matthew, Ph.D Carp, Joshua, Francis J Tinney, Ph.D Bingham, C. Raymond, Ph.D Shope, Jean T., Ph.D Ouimet, Marie Claude, Ph.D Pradhan, Anuj K., and Ed.D Simons-Morton, Bruce G. 2014. Neural Responses to Exclusion Predict Susceptibility to Social Influence. Journal of adolescent health 54, 5 (2014), S22–s31.
- Fardouly et al. (2017) Jasmine Fardouly, Rebecca T. Pinkus, and Lenny R. Vartanian. 2017. The impact of appearance comparisons made through social media, traditional media, and in person in women’s everyday lives. Body image 20 (2017), 31–39.
- Forgas (2011) Joseph P Forgas. 2011. Affective influences on self-disclosure: Mood effects on the intimacy and reciprocity of disclosing personal information. Journal of personality and social psychology 100, 3 (2011), 449.
- Frey et al. (2018) Jérémy Frey, May Grabli, Ronit Slyper, and Jessica Cauchard. 2018. Breeze: Sharing Biofeedback through Wearable Technologies. In Proceedings of the 2018 CHI Conference on human factors in computing systems (Chi ’18). Acm, Ithaca, 1–12.
- Gardner (2013) Howard Gardner. 2013. The app generation : how today’s youth navigate identity, intimacy, and imagination in a digital world.
- Gervais et al. (2016) Renaud Gervais, Jérémy Frey, Alexis Gay, Fabien Lotte, and Martin Hachet. 2016. TOBE: Tangible Out-of-Body Experience. In Proceedings of the TEI ’16: Tenth International Conference on tangible, embedded, and embodied interaction (Tei ’16). Acm, Ithaca, 227–235.
- Goffman (2016) Erving Goffman. 2016. The presentation of self in everyday life. In Social Theory Re-Wired. Routledge, 482–493.
- Greenwald and Breckler (1985) Anthony G Greenwald and Steven J Breckler. 1985. To whom is the self presented. The self and social life 126 (1985), 145.
- Haimson et al. (2021) Oliver L Haimson, Tianxiao Liu, Ben Zefeng Zhang, and Shanley Corvite. 2021. The Online Authenticity Paradox: What Being “Authentic” on Social Media Means, and Barriers to Achieving It. Proc. ACM Hum.-Comput. Interact. 5, Cscw2 (Oct. 2021), 1–18.
- Hamilton et al. (2022) Jessica L Hamilton, Jacqueline Nesi, and Sophia Choukas-Bradley. 2022. Reexamining Social Media and Socioemotional Well-Being Among Adolescents Through the Lens of the COVID-19 Pandemic: A Theoretical Review and Directions for Future Research. Perspect. Psychol. Sci. 17, 3 (May 2022), 662–679.
- Hargittai and Walejko (2008) Eszter Hargittai and Gina Walejko. 2008. The participation divide: Content creation and sharing in the digital age. Information, Community and Society 11, 2 (2008), 239–256.
- Hogan (2010) Bernie Hogan. 2010. The Presentation of Self in the Age of Social Media: Distinguishing Performances and Exhibitions Online. Bull. Sci. Technol. Soc. 30, 6 (2010), 377–386.
- Holt-Lunstad et al. (2010) Julianne Holt-Lunstad, Timothy B Smith, and J Bradley Layton. 2010. Social relationships and mortality risk: a meta-analytic review. PLoS medicine 7, 7 (2010), e1000316.
- House et al. (1988) James S House, Karl R Landis, and Debra Umberson. 1988. Social relationships and health. Science 241, 4865 (1988), 540–545.
- Howell et al. (2019) Noura Howell, Greg Niemeyer, and Kimiko Ryokai. 2019. Life-Affirming Biosensing in Public: Sounding Heartbeats on a Red Bench. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems (Glasgow, Scotland Uk) (Chi ’19, Paper 680). Association for Computing Machinery, New York, NY, USA, 1–16.
- Hu (2022) Cindy Hsinyu Hu. 2022. Self-Disclosure for Early Relationship Development through Situated Prompts in Opportunistic Collective Experiences. In CHI Conference on Human Factors in Computing Systems Extended Abstracts. Acm, New Orleans LA USA, 1–5.
- Huang and Vitak (2022) Xiaoyun Huang and Jessica Vitak. 2022. “Finsta gets all my bad pictures”: Instagram Users’ Self-Presentation Across Finsta and Rinsta Accounts. Proc. ACM Hum.-Comput. Interact. 6, Cscw1 (March 2022), 1–25.
- Impett et al. (2008) Emily A Impett, Lynn Sorsoli, Deborah Schooler, James M Henson, and Deborah L Tolman. 2008. Girls’ relationship authenticity and self-esteem across adolescence. Dev. Psychol. 44, 3 (2008), 722–733.
- Ito (2013) Mizuko Ito. 2013. Hanging out, messing around, and geeking out: Kids living and learning with new media. The MIT press.
- Johnson (1981) John A Johnson. 1981. The” self-disclosure” and” self-presentation” views of item response dynamics and personality scale validity. Journal of Personality and Social Psychology 40, 4 (1981), 761.
- Jones (1990) Edward E Jones. 1990. Interpersonal perception. WH Freeman/Times Books/Henry Holt & Co.
- Kelly et al. (2017) Ryan Kelly, Daniel Gooch, Bhagyashree Patil, and Leon Watts. 2017. Demanding by Design: Supporting Effortful Communication Practices in Close Personal Relationships. In Proceedings of the 2017 ACM Conference on Computer Supported Cooperative Work and Social Computing. Acm, Portland Oregon USA, 70–83.
- Kreling et al. (2021) Rebekka Kreling, Adrian Meier, and Leonard Reinecke. 2021. Feeling authentic on social media: Subjective authenticity across Instagram Stories and Posts. preprint. PsyArXiv. https://doi.org/10.31234/osf.io/jz3wm
- Laurenceau et al. (1998) J Laurenceau, L F Barrett, and P Pietromonaco. 1998. Intimacy as an interpersonal process: the importance of self-disclosure, partner disclosure, and perceived partner responsiveness in interpersonal exchanges. J. Pers. Soc. Psychol. 74, 5 (May 1998), 1238–1251.
- Leary (2019) Mark R Leary. 2019. Self-presentation: Impression management and interpersonal behavior. Routledge.
- Litt (2012) Eden Litt. 2012. Knock, Knock. Who’s There? The Imagined Audience. J. Broadcast. Electron. Media 56, 3 (July 2012), 330–345.
- Liu et al. (2019) Fannie Liu, Mario Esparza, Maria Pavlovskaia, Geoff Kaufman, Laura Dabbish, and Andrés Monroy-Hernández. 2019. Animo: Sharing Biosignals on a Smartwatch for Lightweight Social Connection. Proceedings of ACM on interactive, mobile, wearable and ubiquitous technologies 3, 1 (2019), 1–19.
- Liu et al. (2021) Fannie Liu, Chunjong Park, Yu Jiang Tham, Tsung-Yu Tsai, Laura Dabbish, Geoff Kaufman, and Andrés Monroy-Hernández. 2021. Significant Otter: Understanding the Role of Biosignals in Communication. In arXiv.org. Cornell University Library, arXiv.org, Ithaca.
- Maddox (2023) Jessica Maddox. 2023. More real, or just more surveillance? Panopticism and shifting authenticity paradigms in BeReal. Convergence 29, 5 (Oct. 2023), 1183–1198.
- Marwick and Boyd (2011) Alice E Marwick and Danah Boyd. 2011. I tweet honestly, I tweet passionately: Twitter users, context collapse, and the imagined audience. New Media & Society 13, 1 (2011), 114–133.
- McPherson et al. (2006) Miller McPherson, Lynn Smith-Lovin, and Matthew E Brashears. 2006. Social isolation in America: Changes in core discussion networks over two decades. American sociological review 71, 3 (2006), 353–375.
- Mentis et al. (2014) Helena M Mentis, Jarmo Laaksolahti, and Kristina Höök. 2014. My Self and You: Tension in Bodily Sharing of Experience. ACM transactions on computer-human interaction 21, 4 (2014), 1–26.
- Mitchell and Boer (2017) Robb Mitchell and Laurens Boer. 2017. Move Closer: Towards Design Patterns To Support Initiating Social Encounters. In Proceedings of the 2017 CHI Conference Extended Abstracts on Human Factors in Computing Systems (Denver, Colorado, USA) (Chi Ea ’17). Association for Computing Machinery, New York, NY, USA, 2781–2787.
- Naegele and Goffman (1956) Kaspar D Naegele and Erving Goffman. 1956. The presentation of self in everyday life. Am. Sociol. Rev. 21, 5 (Oct. 1956), 631.
- Peake et al. (2013) Shannon J. Peake, Thomas J. Dishion, Elizabeth A. Stormshak, William E. Moore, and Jennifer H. Pfeifer. 2013. Risk-taking and social exclusion in adolescence: Neural mechanisms underlying peer influences on decision-making. NeuroImage (Orlando, Fla.) 82 (2013), 23–34.
- Peets and Hodges (2018) Kätlin Peets and Ernest V E Hodges. 2018. Authenticity in friendships and well-being in adolescence. Soc. Dev. 27, 1 (2018), 140–153.
- Phua et al. (2017) Joe Phua, Seunga Venus Jin, and Jihoon (Jay) Kim. 2017. Uses and gratifications of social networking sites for bridging and bonding social capital: A comparison of Facebook, Twitter, Instagram, and Snapchat. Computers in Human Behavior 72 (July 2017), 115–122. https://doi.org/10.1016/j.chb.2017.02.041
- Pitt et al. (2021) Caroline Pitt, Ari Hock, Leila Zelnick, and Katie Davis. 2021. The kids are/not/sort of all right. In Proceedings of the 2021 CHI conference on human factors in computing systems. 1–14.
- Piwek and Joinson (2016) Lukasz Piwek and Adam Joinson. 2016. “What do they snapchat about?” Patterns of use in time-limited instant messaging service. Computers in Human Behavior 54 (Jan. 2016), 358–367. https://doi.org/10.1016/j.chb.2015.08.026
- Prinstein (2023) Mitch Prinstein. 2023. Written Testimony. Technical Report. American Psychological Association (APA).
- Qin et al. (2021) Chao Ying Qin, Jun-Ho Choi, Marios Constantinides, Luca Maria Aiello, and Daniele Quercia. 2021. Having a Heart Time? A Wearable-based Biofeedback System. In 22nd International Conference on Human-Computer Interaction with Mobile Devices and Services (Oldenburg, Germany) (MobileHCI ’20, Article 29). Association for Computing Machinery, New York, NY, USA, 1–4.
- Reddy and Kumar (2024) Ananya Reddy and Priya C Kumar. 2024. ‘A Teaspoon of Authenticity’: Exploring How Young Adults BeReal on Social Media. In Proceedings of the CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24, Article 907). Association for Computing Machinery, New York, NY, USA, 1–14. https://doi.org/10.1145/3613904.3642690
- Robinson et al. (2020) Raquel Breejon Robinson, Elizabeth Reid, James Collin Fey, Ansgar E Depping, Katherine Isbister, and Regan L Mandryk. 2020. Designing and Evaluating ’In the Same Boat’, A Game of Embodied Synchronization for Enhancing Social Play. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (Chi ’20). Association for Computing Machinery, New York, NY, USA, 1–14.
- Ryan and Ryan (2019) William S. Ryan and Richard M. Ryan. 2019. Toward a Social Psychology of Authenticity: Exploring Within-Person Variation in Autonomy, Congruence, and Genuineness Using Self-Determination Theory. Review of General Psychology 23, 1 (March 2019), 99–112. https://doi.org/10.1037/gpr0000162
- Schaadhardt et al. (2023) Anastasia Schaadhardt, Cory Gennari Pratt, Yue Fu, and Wanda Pratt. 2023. “Laughing so I don’t cry”: How TikTok users employ humor and compassion to connect around psychiatric hospitalization. (2023).
- Schlenker (1980) Barry R Schlenker. 1980. Impression management: The self-concept, social identity, and interpersonal relations. Social Identity, and Interpersonal Relations (Brooks/Cole, Monterey, CA) (1980).
- Schlenker (1985) Barry R Schlenker. 1985. Identity and self-identification. The self and social life 65, 1 (1985), 99–106.
- Schlenker (2003) Barry R Schlenker. 2003. Self-presentation. Handbook of self and identity 2 (2003), 542–570.
- Schlenker et al. (1996) Barry R Schlenker, Thomas W Britt, and John Pennington. 1996. Impression regulation and management: Highlights of a theory of self-identification. (1996).
- Schlenker and Pontari (2000) Barry R Schlenker and Beth A Pontari. 2000. The strategic control of information: Impression management and self-presentation in daily life. (2000).
- Schlenker and Wowra (2003) Barry R Schlenker and Scott A Wowra. 2003. Carryover effects of feeling socially transparent or impenetrable on strategic self-presentation. Journal of Personality and Social Psychology 85, 5 (2003), 871.
- Schlosser (2020) Ann E Schlosser. 2020. Self-disclosure versus self-presentation on social media. Curr Opin Psychol 31 (Feb. 2020), 1–6.
- Snyder (2024) Sarah J Snyder. 2024. Always-on authenticity: Challenging the BeReal ideal of “being real”. Media Cult. Soc. 46, 2 (March 2024), 404–413.
- Somerville (2013a) Leah H Somerville. 2013a. Special issue on the teenage brain: Sensitivity to social evaluation. Curr. Dir. Psychol. Sci. 22, 2 (April 2013), 121–127.
- Somerville (2013b) Leah H. Somerville. 2013b. The Teenage Brain: Sensitivity to Social Evaluation. Current directions in psychological science : a journal of the American Psychological Society 22, 2 (2013), 121–127.
- Steinberg (2005) Laurence Steinberg. 2005. Cognitive and affective development in adolescence. Trends Cogn. Sci. 9, 2 (Feb. 2005), 69–74.
- Stepanova et al. (2022) Ekaterina R Stepanova, John Desnoyers-Stewart, Kristina Höök, and Bernhard E Riecke. 2022. Strategies for Fostering a Genuine Feeling of Connection in Technologically Mediated Systems. In CHI Conference on Human Factors in Computing Systems. Acm, New Orleans LA USA, 1–26.
- Stephanie Rivas-Lara Becky Pham Jasmine Baten Annie Meyers Yalda T. Uhls (2022) Stephanie Rivas-Lara Becky Pham Jasmine Baten Annie Meyers Yalda T. Uhls. 2022. CSS TEENS & SCREENS 2022 #Authenticity. Technical Report Center for Scholars and Storytellers. Ucla.
- Taber and Whittaker (2020) Lee Taber and Steve Whittaker. 2020. “On Finsta, I can say ’Hail Satan”’: Being Authentic but Disagreeable on Instagram. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems. Acm, Honolulu HI USA, 1–14.
- Tao and Ellison (2023) Michelle Tao and Nicole B Ellison. 2023. “It’s Your Finsta at the End of the Day . . . Kind of”: Understanding Emerging Adults’ Self-Presentational Changes on Secondary Accounts. Social Media + Society 9, 1 (Jan. 2023), 20563051231152812.
- Taylor (2023) Zari A Taylor. 2023. Everyone Stop What You’re Doing and BeReal: Live Networked Publics and Authenticity on BeReal. Social Media + Society 9, 4 (Oct. 2023), 20563051231216959.
- Tetlock and Manstead (1985) Philip E Tetlock and Antony S Manstead. 1985. Impression management versus intrapsychic explanations in social psychology: A useful dichotomy? Psychological review 92, 1 (1985), 59.
- Tice et al. (1995) Dianne M Tice, Jennifer L Butler, Mark B Muraven, and Arlene M Stillwell. 1995. When modesty prevails: Differential favorability of self-presentation to friends and strangers. Journal of personality and social psychology 69, 6 (1995), 1120.
- Tidwell and Walther (2002) Lisa Collins Tidwell and Joseph B Walther. 2002. Computer-mediated communication effects on disclosure, impressions, and interpersonal evaluations: Getting to know one another a bit at a time. Human communication research 28, 3 (2002), 317–348.
- Tirocchi (2023) Simona Tirocchi. 2023. Generation Z, values, and media: from influencers to BeReal, between visibility and authenticity. Front Sociol 8 (2023), 1304093.
- Treem and Leonardi (2013) Jeffrey W Treem and Paul M Leonardi. 2013. Social Media Use in Organizations: Exploring the Affordances of Visibility, Editability, Persistence, and Association. Annals of the International Communication Association 36, 1 (2013), 143–189.
- Tyler (2012) James M Tyler. 2012. Triggering self-presentation efforts outside of people’s conscious awareness. Personality and Social Psychology Bulletin 38, 5 (2012), 619–627.
- Ullman (1987) Chana Ullman. 1987. From Sincerity to Authenticity: Adolescents’ Views of the “True Self”. J. Pers. 55, 4 (1987), 583–595.
- Vanhoffelen et al. (2023) Gaëlle Vanhoffelen, Lara Schreurs, Anneleen Meeus, Nele Janssens, Kathleen Beullens, and Laura Vandenbosch. 2023. BeReal, Be Happy? Examining the relationships between authentic self-presentations on BeReal and adolescents’ self-esteem. New Media & Society (Nov. 2023), 14614448231207783.
- Vaterlaus et al. (2016) J. Mitchell Vaterlaus, Kathryn Barnett, Cesia Roche, and Jimmy A. Young. 2016. “Snapchat is more personal”: An exploratory study on Snapchat behaviors and young adult interpersonal relationships. Computers in Human Behavior 62 (Sept. 2016), 594–601. https://doi.org/10.1016/j.chb.2016.04.029
- Vitak (2012) Jessica Vitak. 2012. The Impact of Context Collapse and Privacy on Social Network Site Disclosures. J. Broadcast. Electron. Media 56, 4 (2012), 451–470.
- Vogels et al. (2022) Emily A Vogels, Risa Gelles-Watnick, and Navid Massarat. 2022. Teens, Social Media and Technology 2022. https://www.pewresearch.org/internet/2022/08/10/teens-social-media-and-technology-2022/. Accessed: 2023-7-7.
- Vohs et al. (2005) Kathleen D Vohs, Roy F Baumeister, and Natalie J Ciarocco. 2005. Self-regulation and self-presentation: regulatory resource depletion impairs impression management and effortful self-presentation depletes regulatory resources. Journal of personality and social psychology 88, 4 (2005), 632.
- Waldinger and Schulz (2016) Robert J Waldinger and Marc S Schulz. 2016. The long reach of nurturing family environments: Links with midlife emotion-regulatory styles and late-life security in intimate relationships. Psychological science 27, 11 (2016), 1443–1450.
- Weinstein and Davis (2017) Emily Weinstein and Katie Davis. 2017. Identity Development in the Digital Age: An Eriksonian Perspective. In Identity, Sexuality, and Relationships among Emerging Adults in the Digital Age. IGI Global, 1–17.
- Williams and Gilovich (2008) Elanor F Williams and Thomas Gilovich. 2008. Do people really believe they are above average? Journal of Experimental Social Psychology 44, 4 (2008), 1121–1128.
- Xiao et al. (2020) Sijia Xiao, Danaë Metaxa, Joon Sung Park, Karrie Karahalios, and Niloufar Salehi. 2020. Random, Messy, Funny, Raw: Finstas as Intimate Reconfigurations of Social Media. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems. Acm, Honolulu HI USA, 1–13.
- Xu et al. (2016) Bin Xu, Pamara Chang, Christopher L. Welker, Natalya N. Bazarova, and Dan Cosley. 2016. Automatic Archiving versus Default Deletion: What Snapchat Tells Us About Ephemerality in Design. In Proceedings of the 19th ACM Conference on Computer-Supported Cooperative Work & Social Computing. Acm, San Francisco California USA, 1662–1675. https://doi.org/10.1145/2818048.2819948
- Yau and Reich (2019) Joanna C Yau and Stephanie M Reich. 2019. “It’s Just a Lot of Work”: Adolescents’ Self-Presentation Norms and Practices on Facebook and Instagram. J. Res. Adolesc. 29, 1 (2019), 196–209.
- Zhao et al. (2008) Shanyang Zhao, Sherri Grasmuck, and Jason Martin. 2008. Identity Construction on Facebook: Digital Empowerment in Anchored Relationships. Computers in Human Behavior 24 (09 2008), 1816–1836. https://doi.org/10.1016/j.chb.2008.02.012
- Zheng et al. (2019) Dong Zheng, Xiao-Li Ni, and Yi-Jun Luo. 2019. Selfie Posting on Social Networking Sites and Female Adolescents’ Self-Objectification: The Moderating Role of Imaginary Audience Ideation. Sex Roles 80, 5 (March 2019), 325–331.
- Zheng et al. (2021) Qingxiao Zheng, Daniela M Markazi, Yiliu Tang, and Yun Huang. 2021. “PocketBot Is Like a Knock-On-the-Door!”: Designing a Chatbot to Support Long-Distance Relationships. Proc. ACM Hum.-Comput. Interact. 5, Cscw2 (Oct. 2021), 1–28.
Appendix A Interview Protocol
<0 Greetings>(5)
Thank you so much for joining me today! Hi, I’m xx, I am a researcher from xx, and I am studying what adolescents and young adults mean by “being real” on social media.
Today we’re going to talk about your experiences with using BeReal, and what you like or dislike about it.
The interview will last no longer than an hour, and we will make sure that none of the information that you share in this interview will be shared with anyone in a personally identifiable way.
Please remember that you should feel free to decline to answer a question if you feel uncomfortable with it.
I wanted to take a moment to ask if you have read the consent agreement and if you have any questions about the consent form?
Great. If you have any questions or concerns regarding this interview during or after our interview, please feel free to email me through the contact information on the consent form.
I don’t work for BeReal, so you don’t need to try to advocate for BeReal if you don’t.
Before we start our interview, do I have your permission to record this interview? The audio may be used in a research paper, but the video will not.
Stop interview.
Great, I’ve just started recording. Today is [date] and I’m talking to [participant ID].
Can you confirm that you have agreed to participate and to be recorded?
<1 Rapport / Apprehension>(10)
How long have you been using BeReal, and how often do(did) you use it?
How did you learn about the platform and why did you decide to use it?
How many friends do you think people typically have on BeReal?
How close are you with your BeReal friends?
How did you select the friends that you added on BeReal?
Do you have friends on BeReal that you don’t really trust?
What’s a typical post that you post(ed) on BeReal? Is there an example that you’re okay with sharing?
How would you describe a typical post that your friends post on BeReal?
Which social media platforms, besides BeReal, do you use or have used before? Which ones did you use most frequently?
What are some of the most fulfilling and the most stressful moments that you experience on each platform?
How is BeReal different from Snapchat, Instagram, and TikTok in terms of its vibe?
What are some biggest differences among each community?
How about Instagram and Snapchat?
What are some reasons that people wouldn’t post on TikTok Now but would post on BeReal?
<2 Exploration>(10)
Do you feel that you’re getting to know your friends better through interactions on BeReal? In what ways?
Which features have you noticed that you think the BeReal creators introduced to encourage people to be more real?
BeReal says that it wants people to be more real online. What’s your impression of that goal?
Do you think BeReal is doing a good job of making a more authentic social media platform? To what extent?
What are some things they can improve on?
Do you ever feel like you’re being forced to be authentic on BeReal because of all the restrictions?
Do you ever hack their system, for example, by waiting on a notification until you’re ready to post, or do you know that your friends do that?
Is it okay or not okay to bend the rules on BeReal? Why?
What are some moments that you experience toxicity on BeReal, if there are any?
Do you ever feel self-conscious about the things you post on BeReal?
If you do, how do you or do you not overcome that?
Have you ever thought that someone else’s post is boring or that someone else’s life seems boring, seeing what they post on BeReal?
What advice would you give to someone who feels self-conscious on BeReal? For example if they are concerned that people will find their posts to be boring or inappropriate?
Was body image issue ever a concern for you or the people you know on BeReal?
Can you give me metaphors?
Do people seem to compare themselves with other people on BeReal? In terms of # of likes, # of friends, what they do on a daily basis, etc.
What’s a cringe post on BeReal? Like that’s weird why would you post that here?
Why do you feel that that is cringy?
Have you heard of any younger/older people that are like 13–15 (16–18) having different views about this? Are they different from the views that friends of your age have?
<3 Co-operative>(15)
Have you ever experienced times when you felt differently about other social media platforms because of your experiences on BeReal?
e.g.) they’re really fake, it feels more wrong to post things there
e.g.) they allow a lot of freedom to post, they are more fun and interesting
Does posting on BeReal feel different from posting on other social media? How?
If there is anything that you would post only on BeReal and not on any other social media, what would it be?
Why is it that on BeReal you can post those things?
What are some things about yourself that are not conveyed well on BeReal?
In your opinion, how would you define authenticity or being real? Why is it important or not important to you?
In your opinion, is authenticity on social media a good thing? Is more authenticity the better?
Are authenticity online and authenticity in daily lives different?
Do you want to experience more authenticity online than what you’re currently experiencing (on BeReal)?
In what context do you feel like you’re the most authentic? At home? On TikTok? At school? With friends?
Can you give an example of a time when you felt like you were being true to yourself?
(if the answer above is offline) What about on online platforms?
Can you think of any backlash or rejection for being authentic on social media?
Do you ever feel any pressure to conform to societal expectations on social media? Does that ever conflict with your efforts to be yourself on social media?
How do you stay authentic to your values and beliefs when faced with peer pressure or temptation?
What’s your impression of the company?
Let’s say you can change current social media in any way possible, from the audience, the design, the features, and everything. What would you like to change the most and why?
When you’re trying to be authentic online, do you find photos to be more effective or text? Temporary vs permanent posts? Larger vs closer groups of friends?
How does BeReal handle privacy and security compared to other platforms you use?
How does BeReal’s content moderation compare to other platforms you use?
Does that affect your view of the platform or the company?
Will having advertisements on BeReal affect how you like BeReal?
<4 Concluding>(5)
How would you describe your BeReal experience using three adjectives? Why did you choose [words]?
If you could change two things about BeReal, what would they be? The design, the features, the people, anything you can think of.
What would make you stop using BeReal? When would you decide to stop using it?